### v1

# Position: Many generalization measures for deep learning are fragile

Abstract
A wide variety of generalization measures have been applied to deep neural networks (DNNs). Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends. In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behaviour. For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalisation measures like the path norm. We also identify subtler forms of fragility. For instance, the PAC-Bayes origin measure is regarded as one of the most reliable, and is indeed less sensitive to hyperparameter tweaks than many other measures. However, it completely fails to capture differences in data complexity across learning curves. This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behaviour, in learning curves, but which is not a post-mortem measure. Beyond demonstrating that many bounds - such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates- are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
Machine Learning, ICML
1 Introduction
Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning. They were developed with two intertwined aims: to furnish tight mathematical guarantees and to explain generalization. For deep networks the same aspiration has driven a profusion of bounds. While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behaviour of deep neural networks (DNNs). These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds. Large-scale comparisons have compared the behaviour of these measures, albeit often only at the sign level, e.g. does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
A notable property of DNNs is that to first order, their generalization performance is robust to modest changes in network size, hyperparameters, optimisers, or stopping criteria. This raises a natural question: are generalization measures themselves equally robust to such changes? If a generalization measure captures the basic underlying factors that drive DNN performance, one would expect it to exhibit a similar degree of robustness.
The main finding of this paper is that many of the post-mortem bounds studied in the literature are fragile: small, ostensibly irrelevant training tweaks that barely affect the underlying DNN, can flip a measure’s shape, trend, or scale. To test these fragilities systematically, we evaluate measures under controlled perturbations that hold data and architecture, fixed while nudging a single training knob. Three stressors structure the audit in the main text: learning‑curve behavior as sample size increases (§3); temporal behavior, including when the model has interpolated (§4); and responses to data complexity via label noise and dataset swaps (§5). In the supplementary material, we show more tests of fragility for a much wider set of measures.
In contrast to the post-mortem bounds, we show that a function-based marginal-likelihood PAC-Bayes bound successfully tracks dataset complexity and learning-curve scaling (Valle-Pérez & Louis, 2020). This bound is based on a simple measure of inductive bias, the probability that a random set of weights produces a function that fits the training data set. Because it operates in function space and does not depend explicitly on training dynamics, this bound is inherently insensitive to the choice of optimiser or most training hyperparameters. This insensitivity is both a strength, making the bound robust rather than fragile, and a limitation, as it prevents the bound from reflecting how training procedures influence generalization, including effects such as feature learning. Nonetheless, the strong performance of this relatively simple approach suggests that developing more sophisticated function-space generalization measures could be a fruitful direction for future research.
Finally, we sketch a few potential causes of fragility, in the hope they can lead to new insights. We show that some of the qualitative failures of norm-based bounds with sample size appear to track a recent prediction of qualitative changes in the scaling of norms in the simpler setting of linear regression (Zhang & Louis, 2025). We also exploit a symmetry in scale‑invariant networks and prove a non‑asymptotic equivalence between fixed learning‑rate with fixed weight decay and a matched run with exponentially increasing learning rate and time‑varying weight decay; the two training procedures compute the same predictor at every iterate (§6; cf. Li & Arora, 2019). Under this invariance lever, magnitude‑sensitive post‑mortem measures can inflate by orders of magnitude while test error remains flat. This suggests that generalization measures should carefully track key invariances present in the underlying DNNs.
The widespread fragility we observe suggests that some generalisation measures may be showing correlation in some limited settings that may not be related to causation. Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
Our contribution:
  • We formulate a compact fragility audit for generalization measures that targets training-hyperparameter stability, post‑interpolation temporal behavior, and data‑complexity response.
  • We provide systematic evidence that popular post‑mortem measures—path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates—change their qualitative predictions under mild optimiser or hyperparameter tweaks even when the underlying DNN accuracy is stable.
  • We prove an equivalence for scale‑invariant networks that matches fixed schedules to exponentially increasing learning‑rate schedules with time‑varying weight decay, yielding a controlled invariance lever that isolates function from parameter scale and exposes magnitude sensitivity.
2 Related work
We use generalization bound and generalization measure interchangeably: both seek to predict out-of-sample performance, differing mainly in whether they arrive with formal guarantees. In modern deep networks these quantities are best read as measures—diagnostics to compare across training/data regimes—rather than practically tight certificates.
From classical capacity to modern practice.
Early theory framed capacity via VC and Rademacher analyses (Mendelson, 2002; Anthony & Bartlett, 2002; Bartlett & Mendelson, 2001; Koltchinskii, 2001). As overparameterization became the norm, evidence accumulated that uniform-convergence explanations often miss the mark: networks can fit random labels yet generalize on natural data (Zhang et al., 2016), some bounds fail to track risk (Nagarajan & Kolter, 2019), and benign overfitting can arise even at interpolation (Bartlett et al., 2020). The field responded with diagnostics that foreground invariances, algorithmic dependence, and data interactions.
2.1 Capacity-oriented diagnostics: norms, margins, distance from initialization
The first works in this subfield translated classical notions into deep settings. Spectrally-normalized margin bounds tied test error to Lipschitz-like control and margins, making scale explicit (Bartlett et al., 2017). Path- and norm-based views clarified how depth and weight scales shape effective complexity (Neyshabur et al., 2015, 2019). Empirically, modeling the distribution of margins—not just the minimum—improves predictiveness across trained families (Jiang et al., 2018; Novak et al., 2018), and sample complexity can depend on norms rather than width (Golowich et al., 2018). In practice, enforcing Lipschitz continuity often improves out-of-sample performance (Gouk et al., 2020; Yoshida & Miyato, 2017). Distance-from-initialization and movement-from-pretraining offer reference-aware, width-robust complexity surrogates with both empirical support and bounds (Li et al., 2018; Zhou et al., 2021; Neyshabur et al., 2017a), while optimization dynamics link large margins to implicit bias in separable regimes (Soudry et al., 2018).
2.2 Geometry-oriented diagnostics: flatness and sharpness
A parallel line approached generalization through loss-landscape geometry. The “flat minima generalize better” intuition predates deep learning (Hochreiter & Schmidhuber, 1997) and resurfaced when large-batch training was observed to converge to sharper minima with worse test error (Keskar et al., 2017). Because raw sharpness is parameterization-dependent, normalized and magnitude-aware definitions were proposed and shown to correlate more robustly with test error (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022). These ideas also shaped algorithms: SAM explicitly optimizes a local worst-case neighborhood and typically improves accuracy; adaptive variants and ablations probe when and why it helps (Foret et al., 2020; Kwon et al., 2021; Andriushchenko & Flammarion, 2022). Independent probes—weight averaging, landscape visualizations, and training-dynamics analyses—add evidence that broader valleys often accompany better generalization (Izmailov et al., 2018; Li et al., 2017; Cohen et al., 2021). Stochastic optimization theory offers a mechanism: noise scale and heavy-tailed gradient noise can bias learning toward flatter regions (Smith et al., 2017; Jastrzebski et al., 2017; Simsekli et al., 2019; McCandlish et al., 2018).
2.3 Algorithm-aware certificates: PAC-Bayes as operational measures
Partly in response to limits of uniform convergence, PAC-Bayesian analysis made algorithm dependence explicit by trading empirical risk against a posterior–prior divergence (McAllester, 1999; Seeger, 2002; Catoni, 2007; Langford & Shawe-Taylor, 2002). The framework became operational when non-vacuous deep-network certificates were obtained by optimizing stochastic posteriors (Dziugaite & Roy, 2017), or simply working with the GP prior (Valle-Pérez et al., 2018; Zhang et al., 2021). Subsequent work tightened certificates with data-aware priors (Dziugaite, 2018; Pérez‐Ortiz et al., 2020), and with compression-flavored posteriors that reduce effective description length (Arora et al., 2018a; Lotfi et al., 2022). Treating PAC-Bayes itself as a training objective yields models that are both accurate and tightly certified, while extensions broaden the scope to adversarial risk and fast/mixed-rate regimes (Rodríguez-Gálvez et al., 2024). However, one worry is that while these optimised measures may provide quantitatively tighter bounds, they do so at the expense of qualitative insight (Arora et al., 2018a).
Other perspectives.
Complementary lenses provide boundary conditions any credible measure should respect. Algorithmic stability ties generalization to insensitivity under data perturbations (Bousquet & Elisseeff, 2002; Hardt et al., 2015); information-theoretic analyses bound excess risk by the information a learning rule extracts from the sample (Russo & Zou, 2016; Xu & Raginsky, 2017). Description-length and compression viewpoints link compressibility to generalization (Blier & Ollivier, 2018; Taiji Suzuki, 2020). Linearized regimes (NTK/GP) delineate when very wide nets behave like their kernel limits (Jacot et al., 2018; Lee et al., 2020b). Finally, double-descent phenomena and scaling laws offer external checks on diagnostics (Belkin et al., 2019; Kaplan et al., 2020).
Meta-evaluations and synthesis.
Large-scale comparisons underscore that no single candidate explains generalization under all interventions (Jiang et al., 2019b; Dziugaite et al., 2020b), motivating a measures-as-diagnostics mindset. A recent critique shifts the lens from predictiveness to tightness, showing that uniformly tight bounds are out of reach in overparameterized settings and clarifying what such tightness could mean (Gastpar et al., 2023). Our focus is complementary: rather than tightness per se, we study fragility—how otherwise informative measures can fail under innocuous hyperparameter changes that leave underlying DNN performance essentially unchanged.
3 Training-hyperparameter fragility
The path norm (definition in App. A) is often regarded as the strongest norm-based proxy for modern ReLU networks: it respects layerwise rescalings and has supported some of the sharpest norm-style capacity statements and practical diagnostics (Neyshabur et al., 2015, 2017b; Gonon et al., 2023). Precisely because it is a strong candidate, it is a natural place to look for fragility.
In Fig. 1 we show a triptych for ResNet‑50 on FashionMNIST, nudging exactly one hyperparameter at a time, either the optimizer or the learning rate (no early stopping), while keeping the data, model, and pipeline fixed. Each curve shows test error (black), the raw path norm (blue), and the path norm divided by the empirical margin (orange); axes are logarithmic and error bars reflect seed variability.
Figure 1: Path‑norm learning curves under small training‑pipeline changes (no early stopping). Left: SGD with momentum, LR =0.01=0.01. Middle: Adam, LR =0.01=0.01. Right: Adam, LR =0.001=0.001. Curves show ResNet‑50 test error (black), path norm (blue), and path norm over margin (orange); axes are logarithmic and error bars denote seed variability. Training set sizes $n$ range from $10^{2}$ to the full FashionMNIST training set ($5\times 10^{4}$ samples).
In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower. Both drop faster than the underlying ResNet‑50 test error. Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behaviour. A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows. These widely different behaviours arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
The contrasting behaviour of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression. In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer. More precisely, with $p\in[1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if $r\leq 2(p-1)$  (Zhang & Louis, 2025). This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm. This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (Zhang & Louis, 2025). For the more complex setup of path-norm for ResNet‑50 on FashinMNIST in Fig. 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
See App. A for precise definitions for the measures we consider as well as further experiments, including the same stress test across Frobenius, margin, spectral, PAC‑Bayes, and VC‑style proxies.
4 Temporal behavior fragility
Temporal traces ask a simple question that post‑mortem snapshots cannot: once the classifier stops changing, does the measure stop moving? Let $T_{\mathrm{int}}$ denote the first epoch at which training accuracy reaches $100\%$. A stable diagnostic should largely settle for $t>T_{\mathrm{int}}$; continued change indicates optimizer‑driven drift rather than functional change.
A set of three experiments on FashionMNIST illustrates the point. We fix the architecture and data, then vary one knob at a time across the three panels of Figure 2. Holding the learning rate fixed at $0.01$ and swapping the optimizer from SGD with momentum (left) to Adam (middle) flips the post‑interpolation regime: with SGD, the path norm sits near $5\times 10^{2}$ and then slowly declines and the ratio‑over‑margin trends downward ($T_{\mathrm{int}}\!\approx\!27$), whereas with Adam the path norm climbs from $\approx 10^{-2}$ to $10^{0}$ and the ratio rises in step, continuing past interpolation ($T_{\mathrm{int}}\!\approx\!114$). Holding the optimizer fixed (Adam) and reducing the learning rate from $0.01$ to $0.001$ (right) again changes the temporal behavior: both traces drop rapidly and then increase moderately after the dashed line. In all cases, the generalization curve remains on a comparable scale.
A plausible mechanism underlies this split. After interpolation under cross‑entropy, gradients can continue to increase logit scale along near‑flat directions. Adam tends to amplify this scale drift, while SGD+momentum tempers it, yielding opposite signs for the post‑interpolation slope. The lesson is that weight–norm–based measures—and margin‑normalized variants that co‑move with them—are not monotone indicators of optimization progress: an upward trend can be optimizer‑driven scale inflation; a flat or gently declining trend need not signal stagnation.
For practice, we recommend reporting $T_{\mathrm{int}}$ and the post‑interpolation slope $s\coloneqq\tfrac{d\log\mathrm{Measure}}{d\log t}\big|_{t>T_{\mathrm{int}}}$ alongside any temporal plot; overlay optimizers on the same axes; and compare snapshots at matched milestones (e.g., the first $100\%$‑accuracy crossing) or with early‑stopped checkpoints. A useful stress test is hysteresis: resume a checkpoint taken just after $T_{\mathrm{int}}$ with a different optimizer and check whether the measure’s drift changes sign while test error holds steady. These checks make temporal fragility visible and keep optimizer‑induced motion from being mistaken for learning. We refer the readers to Appendix B for experimental results on the temporal behaviors of more generalization measures.
Figure 2: ResNet‑50 on FashionMNIST as a function of training epochs. Left (SGD+momentum, LR =0.01=0.01). Middle (Adam, LR =0.01=0.01). Right (Adam, LR =0.001=0.001). While the path norm and path norm over margin measures show qualitative different behavior upon small hyperparameter tweaks, the underlying generalization is similar across the three panels.
5 Data-Complexity Fragility: Label Noise & Dataset Difficulty
How DNN performance varies with data complexity or sample size (learning curves) offers key insights into the mechanisms of generalization. Therefore, reporting how a generalization measure behaves under changes in data complexity or sample size should be a standard part of its evaluation.
The first two panels in figure 3 illustrate how merely changing Adam’s learning rate flips the qualitative trend of the path‑norm curves (left vs. middle) with data complexity (for SGD with momentum on this data, see Appendix D). This illustrates the kind of fragility our framework urges authors to investigate and report when treating bounds as measures.
Finally, the last panel in figure 3 illustrates how the function‑space ML‑PACBayes measure provides a non-vacuous bound that tracks data-complexity (see also Appendix E). This performance is not mirrored by several PAC–Bayes parameter‑space surrogates, which on the one hand are much less sensitive to training hyperparameters than the path norm measures are, but on the other hand exhibit a different more subtle kind of fragility; they are typically insensitive to label corruption in standard setups (see Appendix D). We call this behaviour a fragility because it reflects a serious failing of these measures.
Figure 3: Label‑corruption sweeps with post‑mortem norms vs. function‑space ML‑PACBayes on RESNET-50. Left: Adam, $\mathrm{LR}{=}0.01$. Middle: Adam, $\mathrm{LR}{=}0.001$. Right: ML‑PACBayes; purple, Dashed black: test error; blue: path norm; orange: path norm over margin. A ten‑fold LR change reverses the path‑norm trend (left vs. middle), while ML‑PACBayes remains a smooth, monotone function of corruption (right) and is, by construction, agnostic to optimizer/LR. All panels use 10,000 training samples on binarized FashionMNIST.
To illustrate data-complexity fragility, we chose the PACBAYES_ORIG measure (definition in App. A) because it is thought to be one of the best in its class. We hold the architecture and hyperparameters fixed and use the datasets MNIST, FashionMNIST, and CIFAR10. As illustrated in Figure 4, the dashed generalization–error curves separate cleanly and their slopes decrease with dataset difficulty (MNIST steepest decline, CIFAR‑10 flattest). However, the first two panels of Figure 4 show that for a given set of hyperparameters, the PACBAYES_ORIG measure gives nearly identical numerical values for different datasets and therefore fails to recognize changes in dataset complexity: its learning‑curve slopes are nearly indistinguishable across datasets and reflect changes in the test‑error slopes. By contrast, the third panel (ML‑PACBayes) is both tight (close to the corresponding error curves) and data‑aware: it preserves the MNIST << FashionMNIST << CIFAR‑10 ordering and reflects the different slopes. See Appendix E for the bound’s definition and a fuller discussion of why this function‑space quantity captures dataset difficulty while post-mortem parameter‑space PAC-Bayes surrogates often do not.
There has been substantial recent interest in the study of scaling laws for deep neural networks (DNNs). These laws describe how generalization performance varies as a function of model size, training sample size, or computational cost (see, e.g., (Hestness et al., 2017; Kaplan et al., 2020; Spigler et al., 2019; Hoffmann et al., 2022; Valle-Pérez & Louis, 2020; Bahri et al., 2021; Nam et al., 2024; Goring et al., 2025) and references therein). Here, we observe that the ML-PACBayes bound does quite well at capturing the scaling behavior with the training sample size. However, it should be kept in mind that because this bound is evaluated with a Gaussian process (GP) method, it cannot capture feature-learning, which can play a crucial role in learning-curve scaling for certain datasets (Nam et al., 2024; Goring et al., 2025).
Figure 4: Fixed architecture, varying dataset.. The model is RESNET-50. Left: PACBAYES_ORIG with Adam. Middle: PACBAYES_ORIG with SGD (momentum). Changing from ADAM to SGD with momentum leads to a small increase in the overall scale of the PACBAYES_ORIG learning curves (solid lines). However, for both optimisers, these learning curves don’t distinguish MNIST, FashionMNIST, and CIFAR10, missing both the data‑complexity ordering and the changing slopes visible in the underlying learning curves (dashed lines). Right: ML‑PACBayes (marginal‑likelihood PAC–Bayes) is tight and preserves the correct dataset ordering and reproduces the trends in the slope; for technical reasons, this panel is computed on the binarized versions of the three datasets.
In App. F, we move beyond label corruption and dataset swaps, and also probe symmetry‑preserving vs. signal‑destroying data transforms via pixel permutations. Again, we find that the ML-PACBayes function‑space predictor tracks the expected invariances while several post‑mortem diagnostics do not.
6 Scale-invariant network and exponential learning rate schedule
Normalization layers make many modern networks effectively scale‑invariant, and this symmetry underlies the “exponentially increasing learning‑rate” idea of Li & Arora (2019): in such models, training with fixed learning rate (LR), weight decay (WD), and momentum can be matched—in function space—by a run with an exponentially increasing LR and no WD (plus a small momentum correction across phases). Their analysis shows how LR and WD can be “folded into” an exponential schedule and documents successful experiments under this equivalence. We refer to this multiplicative control parameter $\alpha$ as the Exp++ factor.
We use the same symmetry but a different equivalence. Instead of removing WD, our theorem pairs an exponentially increasing LR with a time‑varying WD so that every iterate computes the same predictor as a fixed‑LR/fixed‑WD baseline. The mapping is explicit and non‑asymptotic: it gives closed‑form schedules for $\tilde{\eta}_{t}$ and $\tilde{\lambda}_{t}$, and an admissible range for the multiplicative factor $\alpha$ via the interval $\mathcal{I}$. Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures. Empirical details and results are deferred to Appendix G.
Definition 6.1 (Scale invariant neural networks).
Consider a parameterized neural network $f(\boldsymbol{\theta})$. We say $f$ is scale invariant if
$\forall c\in\mathbb{R}^{+},f(\boldsymbol{\theta})=f(c\boldsymbol{\theta})$ (1)
Theorem 6.2 (Equivalence of schedules in scale-invariant nets).
[…]
$\Delta_{\lambda}$ $\coloneqq(1-\gamma)^{2}-2(1+\gamma)\lambda\eta_{0}+(\lambda\eta_{0})^{2},$ (2)
$\Xi(\alpha)$ $\coloneqq\frac{\alpha^{2}-\alpha(1-\lambda\eta_{0}+\gamma)+\gamma}{\eta_{0}}.$ (3)
Define the interval endpoints
$\alpha_{L}$ $\coloneqq\frac{\gamma}{\,1-\lambda\eta_{0}+\gamma\,}$, (4)
$\alpha_{-}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}-\sqrt{\Delta_{\lambda}}}{2},$ (5)
$\alpha_{+}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}+\sqrt{\Delta_{\lambda}}}{2},$ (6)
and set
$\mathcal{I}\coloneqq(\alpha_{L},\ \alpha_{-}]\ \cup\ [\alpha_{+},\,1).$ (7)
Consider the two updates (with $\theta_{-1}=\theta_{0}$, $\tilde{\theta}_{0}=\theta_{0}$, $\tilde{\theta}_{-1}=\alpha\,\theta_{-1}$):
$\frac{\theta_{t}-\theta_{t-1}}{\eta_{0}}$ $=\frac{\gamma(\theta_{t-1}-\theta_{t-2})}{\eta_{0}}-\nabla_{\theta}\!\Big(L(\theta_{t-1})+\tfrac{\lambda}{2}\|\theta_{t-1}\|_{2}^{2}\Big),$ (A)
$\frac{\tilde{\theta}_{t}-\tilde{\theta}_{t-1}}{\tilde{\eta}_{t}}$ $=\frac{\gamma(\tilde{\theta}_{t-1}-\tilde{\theta}_{t-2})}{\tilde{\eta}_{t-1}}-\nabla_{\theta}\!\Big(L(\tilde{\theta}_{t-1})+\tfrac{\tilde{\lambda}_{t}}{2}\|\tilde{\theta}_{t-1}\|_{2}^{2}\Big).$ (B)
If $\alpha\in\mathcal{I}$, then the schedules
$\tilde{\eta}_{t}$ $=\eta_{0}\,\alpha^{-2t-1},$ (8)
$\tilde{\lambda}_{t}$ $=\Xi(\alpha)\,\alpha^{2t-1}+\frac{\gamma(1-\alpha)}{\eta_{0}\,\alpha}\,\mathbbm{1}\{t=0\}$ (9)
ensure $\tilde{\theta}_{t}=\alpha^{-t}\theta_{t}$ for all $t\geq 0$; hence (A) and (B) generate identical functions.
Remark 6.3 (Parameter range and nonnegativity).
Assume
$\lambda\eta_{0}$ $\leq(1-\sqrt{\gamma})^{2},$ (10)
$\alpha_{L}$ $<\alpha_{-}\quad\text{(equivalently, $\ \Delta_{\lambda}\geq 0$)}.$ (11)
Then the interval $\mathcal{I}$ above is nonempty and the square‑root term is real; both conditions hold for common hyperparameters in practice.
Consequences. The proof of Theorem 6.2 is in Appendix C. In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures. Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix G.
7 Quantitative fragility: a scale‑free instability score
The heuristic stress tests in Sections 35 show that several popular post‑mortem measures are fragile. Here we provide a complementary quantitative lens: scale‑free statistics that capture how much a measure fluctuates among runs with (nearly) the same test error. These quantitative metrics of fragility can facilitate larger-scale comparisons between measures.
Setup.
Fix a dataset/architecture group $g$ with trained runs $\mathcal{S}_{g}$. For run $r\in\mathcal{S}_{g}$, let $\widehat{\varepsilon}_{r}\in[0,1]$ denote test error and $C_{r}\in(0,\infty)$ the value of a post‑mortem measure $C$ (e.g., a norm proxy). For an absolute error tolerance $\delta>0$, define the close‑error set
$\Pi_{\delta}(\mathcal{S}_{g})\;=\;\bigl\{(r,s)\in\mathcal{S}_{g}^{2}:\ r<s,\ \lvert\widehat{\varepsilon}_{r}-\widehat{\varepsilon}_{s}\rvert\leq\delta\bigr\}.$ (12)
Definition 7.1 (Conditional Measure Spread (CMS)).
The CMS of $C$ in group $g$ at tolerance $\delta$ is
$\mathrm{CMS}_{\delta}(C;g)\;=\;\operatorname{median}_{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g})}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (13)
Interpretation. $e^{\mathrm{CMS}_{\delta}}$ is the typical multiplicative factor by which $C$ varies across runs that are indistinguishable in test error up to $\delta$ (e.g., $\mathrm{CMS}_{\delta}\!\approx\!\ln 2$ means a factor‑two swing).
Randomness adjustment.
To separate hyperparameter induced jumpiness from random‑seed jitter, let $H_{r}$ collect the training hyperparameters for run $r$ (e.g., optimizer, learning rate, schedule flags, early‑stopping rule, training‑set size), and let $\textsf{seed}_{r}$ be its random seed. Split close‑error pairs into
$\Pi_{\delta}^{\text{seed}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}=H_{s},\ \textsf{seed}_{r}\neq\textsf{seed}_{s}\bigr\},$ (14)
$\Pi_{\delta}^{\text{inter}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}\neq H_{s}\bigr\}.$ (15)
Define the corresponding spreads
$\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{seed}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$, (16)
$\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{inter}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (17)
Definition 7.2 (eCMS (Excess CMS)).
The seed‑adjusted excess at matched error is
$\mathrm{eCMS}_{\delta}(C;g)\;=\;\Bigl[\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)-\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)\Bigr]_{+}.$ (18)
Aggregation across groups.
Let $\mathcal{G}$ be the set of dataset/architecture groups. We summarize by medians across groups (and report coverage):
$\mathrm{CMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{CMS}_{\delta}(C;g)\bigr\},$ (19)
$\mathrm{eCMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{eCMS}_{\delta}(C;g)\bigr\}.$ (20)
This “median‑of‑medians” is robust to heterogeneous group sizes and outliers.
How to read the scores.
  • CMS. Values near 0 indicate a stable measure among equal‑error runs; larger values mean greater local instability. (For reference, $\ln 2\approx 0.69$, $\ln 3\approx 1.10$ in the same log units.)
  • eCMS. Values near 0 indicate that most equal‑error variability is seed‑level; large eCMS points to hyperparameter‑sensitive jumpiness beyond seeds.
  • Effect of $\delta$. Decreasing $\delta$ tightens the notion of “equal‑error.” Scores that remain large as $\delta\!\downarrow\!0$ reflect genuinely volatile measures.
Properties and protocol.
Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$). Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (12)–(15) and the fraction of groups covered. Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
8 Conclusion and discussion
We take a pragmatic view of generalization bounds: in modern deep learning they function best as measures to be judged by how they behave across data and training regimes rather than by worst‑case tightness. Read this way, fragility becomes a key property to audit. With the task fixed and accuracy essentially steady, small and reasonable pipeline tweaks—changing the learning rate by a factor of ten, swapping Adam for momentum SGD, toggling early stopping—can bend surrogate curves, invert trends, or inflate their scale. When the task complexity changes, via label corruption or by moving from MNIST to FashionMNIST to CIFAR‑10, several surrogates fail to preserve the expected ordering or slope. Our scale‑invariance construction makes this failure mode explicit: one can inflate parameter magnitudes without changing the predictor, and magnitude‑sensitive post‑mortems then balloon while test error does not. This gap underscores that properties of a single parameter vector need not reflect properties of the learned function.
The comparison between post‑mortem diagnostics and the marginal‑likelihood PAC–Bayes route in function space sharpens the lesson. This predictor can capture key trends with data complexity, including the slopes of learning curves. Given that the interaction between data and architecture is one of the key components of DNN performance that needs explaining, the success of this bound, and its simple interpretation in terms of inductive bias at initialisation, provides a concrete foil against which parameter‑space surrogates can be refined.
Looking forward, several concrete directions could strengthen this position:
  • Design reparameterization‑invariant, data‑aware post‑training diagnostics that borrow the Occam flavour of marginal likelihood while remaining practical at scale;
  • build a public fragility benchmark and automated audit harness covering optimizer/schedule swaps, scale‑symmetry probes, and data‑complexity sweeps.
  • Develop approximate function‑space surrogates (finite‑width corrections, amortized evidence estimators, ensembles as posteriors over functions) to make invariance‑friendly predictors usable in routine training.
  • Tie diagnostics to training dynamics by discouraging scale‑only drift after interpolation or by incorporating function‑space objectives during training.
  • Extend evaluation beyond vision classification to sequence models, generative modelling, reinforcement learning, and explicit distribution shift, testing whether the same invariance principles hold.
9 Alternative Views
Reasonable researchers can disagree with our emphasis on parameter-space fragility. We highlight several viable counterpositions and address them in turn:
  • Optimizer sensitivity is signal, not noise. Because implicit bias drives which predictor is learned, a useful measure should move with the training pipeline, not be invariant to it.
  • Post‑mortem measures can be stabilized. With the right reparameterization‑aware definitions (e.g., normalized sharpness, margin distributions, distance from initialization), much of the apparent fragility disappears.
  • Function‑space GP baselines are mis‑specified for modern, finite‑width practice. GP priors, absence of data augmentation in the prior, and finite‑width effects limit how decisively marginal‑likelihood predictors should be used as a reference.
Optimizer sensitivity as a feature.
It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalisation measures. This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019). From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
Response. We agree that optimizer choice can change the learned function and that a diagnostic should detect functionally meaningful changes. Our claim is narrower: diagnostics that swing under pure scale drift or minor hyperparameter nudges while test error and predictions remain essentially fixed can mislead day‑to‑day comparisons. This is why our audit always pairs any optimizer comparison with either (i) matched‑prediction checkpoints (e.g., at the first $100\%$‑accuracy crossing) or (ii) a symmetry lever that alters parameter scale without changing the predictor (§6). When pipeline changes do alter the predictor, a stable measure should move in tandem with test error; when they do not, a robust measure should be indifferent.
Stabilizing post‑mortem diagnostics.
A second objection holds that many post‑training measures already address scale and parameterization issues. Normalized or reparameterization‑aware sharpness correlates more reliably with generalization than raw curvature (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022); modeling the distribution of margins rather than the minimum improves predictiveness across runs (Jiang et al., 2018); reference‑aware surrogates such as distance from initialization or movement from pretraining reduce width and scale artifacts and come with supporting analyses (Li et al., 2018; Neyshabur et al., 2017a; Zhou et al., 2021). On this view, fragility largely reflects naive implementations, not intrinsic flaws.
Response. We see these developments as complementary and encouraging. Our results target precisely such “best‑effort” variants (normalized, margin‑aware, reference‑aware) and still uncover qualitative flips under mild pipeline changes when predictions are stable. We do not argue that post‑mortems are without merit; rather, we argue they remain fragile enough that authors should routinely report their stability under the kinds of stressors we outline (§35), and prefer versions that (i) are explicitly invariant to layerwise rescalings, (ii) condition on matched prediction milestones, and (iii) are benchmarked against data‑difficulty shifts and temporal drift after interpolation. In short, our audit is a bar to clear, not a dismissal of the enterprise.
Limits of function–space GP references.
A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix E) overstates its practical authority. Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination. While GPs are known to closely match DNN performance on the datasets we test in this paper (Lee et al., 2020b), it is hard to see how one could untangle the effects of hyperparameters or optimiser choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address. More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified. In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys. Therefore, the GP route’s stability could stem from omitting factors that matter in practice (Lee et al., 2020b; Valle-Pérez & Louis, 2020).
Response. We agree that one of the most important questions that needs explaining in DNNs is how, when, and why finite‑width networks outperform their infinite-width GP limits, in particular for settings where feature-learning is important (see e.g. (Goring et al., 2025) and references therein). We use the GP based marginal‑likelihood bound as a calibration tool, not an oracle: it naturally encodes desirable invariances (insensitivity to parameter scale and optimizer path once the dataset is fixed) and for the cases we study, consistently tracks data difficulty and learning‑curve scaling across families. Those properties make it a valuable foil for stress‑testing post‑mortem bounds. Where the GP prior is clearly misspecified (e.g., heavy augmentation or domains where feature-leaning is key), our recommendation is empirical: rerun the same fragility audit. If the GP predictor ceases to track error while a reparameterization‑invariant post‑mortem does, that is evidence for the importance of a post‑mortem analysis in that regime.
Synthesis.
These counterarguments motivate a middle path. Rather than elevating any single diagnostic, we advocate (i) reporting multiple invariance‑aware post‑mortem measures, (ii) auditing them with our stressors (learning‑curve shape, temporal drift after interpolation, and data‑complexity response), and (iii) anchoring interpretation with at least one function–space reference when feasible. This recognizes optimizer‑dependent mechanisms while discouraging conclusions drawn from features (e.g., raw scale) that can be changed without affecting predictions.
Impact Statement
This is a position paper about evaluation practice. The main impact is methodological: a simple audit protocol and reporting guidelines that encourage robustness checks for generalization measures. Potential positive impacts include more reliable model selection and clearer communication of limitations; potential negative impacts are minimal, as we do not propose deployable systems or datasets.
Appendix A Training-hyperparameter fragility for all measures
In this section we give concise definitions of the measures we study (aligning notation with prior large-scale evaluations) and provide additional experimental evidence beyond the main text (Dziugaite et al., 2020b; Jiang et al., 2019b). We then present figure-backed comparisons where small learning-rate or optimizer tweaks trigger qualitatively different curve shapes for the surrogate—plateaus, rebounds, late spikes, and crossings—while the accompanying accuracy curves remain comparatively calm.
A.1 Measures considered
We adopt the categories and normalizations used by prior experimental studies of generalization measures (Jiang et al., 2019b; Dziugaite et al., 2020b); for exact constants and implementation choices, see App. C.6 of Dziugaite et al. (2020b). Let a $d$-layer network have layer weights $\{W_{i}\}_{i=1}^{d}$ with initialization $\{W_{i}^{0}\}$; write $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ for Frobenius and spectral norms; let $n$ denote training-set size and $\gamma$ a robust (e.g., 10th-percentile) training margin.
  • Frobenius distances. Layerwise distances from initialization and norm aggregates, e.g.
    $C_{\mathrm{FrobDist}}\;=\;\sqrt{\sum_{i=1}^{d}\|W_{i}-W_{i}^{0}\|_{F}^{2}\;/\;n}\,,\qquad C_{\mathrm{param}}\;=\;\sqrt{\sum_{i=1}^{d}\|W_{i}\|_{F}^{2}\;/\;n}\,.$
  • Inverse margin. A margin-based surrogate,
    $C_{\mathrm{inv\text{-}margin}}\;\propto\;\frac{\sqrt{n}}{\gamma}\,.$
  • Spectral metrics. Products/means and distances in spectral norm, e.g.
    $C_{\prod\mathrm{spec}}\;=\;\sqrt{\prod_{i=1}^{d}\|W_{i}\|_{2}^{2}\;/\;n}\,,\qquad C_{\mathrm{DistSpecInit}}\;=\;\sqrt{\sum_{i=1}^{d}\|W_{i}-W_{i}^{0}\|_{2}^{2}\;/\;n}\,.$
  • Combined spectral–Frobenius ratio. We follow the combined ratio used in prior large‑scale studies (App. C.6 of Dziugaite et al., 2020a), denoted FRO_OVER_SPEC, which normalizes Frobenius quantities by spectral ones to reduce raw scale effects. We report it alongside its constituents in our plots.
  • PAC-Bayes families and flatness proxies. Bounds/proxies parameterized by posterior radii $\sigma$ (and magnitude-aware $\sigma_{0}$), e.g.
    $C_{\mathrm{PACBAYES\text{-}ORIG}}\;=\;\frac{1}{\sqrt{n}}\sqrt{\frac{\|w\|_{2}^{2}}{4\sigma^{2}}+\log\!\Big(\frac{n}{\delta}\Big)+10},\quad C_{\mathrm{Flatness}}\;=\;\frac{1}{\sigma\sqrt{n}},\quad C_{\mathrm{MAG\text{-}Flatness}}\;=\;\frac{1}{\sigma_{0}\sqrt{n}}.$
  • Path norms. With $w=\mathrm{vec}(W_{1},\ldots,W_{d})$ and $f_{w^{2}}(\mathbf{1})[i]$ the $i$th logit when all weights are squared elementwise and the input is all ones,
    $C_{\mathrm{path.norm}}\;=\;\sqrt{\sum_{i}f_{w^{2}}(\mathbf{1})[i]\;/\;n},\qquad C_{\mathrm{path.norm\text{-}over\text{-}margin}}\;=\;\sqrt{\sum_{i}f_{w^{2}}(\mathbf{1})[i]\;/\;(\gamma^{2}n)}.$
  • VC-dimension proxy (parameter count). A coarse parameter-count surrogate,
    $C_{\mathrm{params}}\;=\;\sqrt{\sum_{i=1}^{d}k_{i}^{2}\,c_{i-1}(c_{i}+1)\;/\;n}\,$,
    with kernel sizes $k_{i}$ and channel counts $c_{i}$.
A.2 Frobenius distance
Two qualitatively different shapes appear when only the learning rate or optimizer changes, and they are visible in Fig. S1 and Fig. S2. On ResNet-50/CIFAR-10 with Adam, $\eta{=}10^{-3}$ produces a broad plateau followed by a late collapse, whereas $\eta{=}10^{-2}$ decays smoothly throughout (compare left vs. right in Fig. S1). On DenseNet-121/FashionMNIST at fixed $\eta{=}10^{-2}$, Adam’s curve is U-shaped (drop then rebound), while momentum SGD traces a near log-linear decline (left vs. right in Fig. S2); the accuracy curves remain closely aligned in both pairs.
Figure S1: Frobenius distance, ResNet-50 on CIFAR-10 (Adam). Left: $\eta{=}10^{-3}$ shows a plateau and late collapse; right: $\eta{=}10^{-2}$ decays smoothly while accuracy tracks similarly.
Figure S2: Frobenius distance, DenseNet-121 on FashionMNIST ($\eta{=}10^{-2}$). Left: Adam yields a U-shape; right: SGD+mom declines nearly log-linearly; in both cases test accuracy evolves similarly.
A.3 Inverse margin
Figure S3 (left vs. right) shows that for ResNet-50/CIFAR-10 with Adam, $\eta{=}10^{-3}$ produces a smooth, power-law-like decline in inverse margin, whereas $\eta{=}10^{-2}$ stalls mid-training before resuming its drop; this kink has no analogue in the accuracy curve. At fixed $\eta{=}10^{-2}$ on FCN/FashionMNIST, Adam develops a clear bump after roughly $10^{3}$ samples while SGD’s curve decreases monotonically (Fig. S4); again, both reach similar generalization.
Figure S3: Inverse margin, ResNet-50 on CIFAR-10 (Adam). Left: $\eta{=}10^{-3}$ decays smoothly; right: $\eta{=}10^{-2}$ stalls then resumes, a kink absent from the accuracy curve.
Figure S4: Inverse margin, FCN on FashionMNIST ($\eta{=}10^{-2}$). Left: Adam develops a mid-course bump; right: SGD+mom is monotone; both generalize similarly.
A.4 Spectral metrics
Spectral surrogates display late spikes and order reversals under the same minimal tweaks. With DenseNet-121/FashionMNIST and Adam, $\eta{=}10^{-3}$ sends the distance-from-initialization in spectral norm down and then sharply up late in training, crossing the FRO-OVER-SPEC curve, whereas $\eta{=}10^{-2}$ keeps both traces monotone but flips their ordering (compare panels in Fig. S5). On ResNet-50/FashionMNIST at $\eta{=}10^{-2}$, Adam shows a valley then rise, while SGD declines steadily (Fig. S6); accuracy curves overlap in both pairs.
Figure S5: Spectral metrics, DenseNet-121 on FashionMNIST (Adam). Left: $\eta{=}10^{-3}$ drops then spikes, crossing FRO-OVER-SPEC; right: $\eta{=}10^{-2}$ stays monotone but reverses ordering.
Figure S6: Spectral metrics, ResNet-50 on FashionMNIST ($\eta{=}10^{-2}$). Left: Adam exhibits a valley then rise; right: SGD+mom declines steadily; accuracy is similar.
A.5 PAC-Bayes bounds
Optimizer swaps and moderate learning-rate changes also produce curved versus straight “bound profiles” and late order crossings. For ResNet-50/FashionMNIST at $\eta{=}10^{-2}$, Adam yields kinked “banana” trajectories across variants, whereas SGD renders near-straight lines (Fig. S7). Holding Adam fixed and raising $\eta$ from $10^{-3}$ to $10^{-2}$ reorders the variants late in training—for example, PACBAYES MAG and PACBAYES ORIG swap rank—even though accuracy shows no such crossing (Fig. S8).
Figure S7: PAC-Bayes, ResNet-50 on FashionMNIST ($\eta{=}10^{-2}$). Left: Adam produces curved, kinked trajectories; right: SGD+mom is nearly straight; test error changes little.
Figure S8: PAC-Bayes, ResNet-50 on FashionMNIST (Adam). Left: $\eta{=}10^{-3}$; right: $\eta{=}10^{-2}$. Increasing $\eta$ induces late-training order swaps (e.g., MAG vs. ORIG) absent from the accuracy curves.
A.6 VC-dimension proxy (robust baseline)
As a control, Fig. S9 shows that the parameter-count proxy for ResNet-50/CIFAR-10 with Adam is strikingly shape-stable across $\eta{=}10^{-3}$ and $\eta{=}10^{-2}$: both traces are near-identical monotone decays, tracking one another while accuracy also aligns. This pair serves as a rare counterexample resistant to qualitative shifts.
Figure S9: VC-dimension proxy, ResNet-50 on CIFAR-10 (Adam). Left: $\eta{=}10^{-3}$; right: $\eta{=}10^{-2}$. Nearly identical monotone decays; a useful control.
Across all families above, the figures make a consistent point: changing only the learning rate or swapping the optimizer can reshape the surrogate’s learning curve—introducing plateaus, rebounds, late spikes, or crossings—without a commensurate shift in test accuracy. This shape-level fragility extends far beyond Path norms and cautions against reading causal explanations of generalization from any single surrogate without dedicated stress tests (Dziugaite et al., 2020b; Jiang et al., 2019b).
Appendix B Additional temporal behavior results: optimizer sensitivity across measure families
This section extends the temporal analysis in Section 4 by holding the task and hyperparameters fixed and changing only the optimizer. We train ResNet–50 on FashionMNIST with learning rate $0.01$ and no early stopping, and we track each measure across epochs. All panels use a logarithmic epoch axis; this makes the early regime and successive orders of magnitude more legible, while very late additive-only epochs occupy little horizontal extent unless they span a substantial multiplicative range. The red dashed vertical line in every panel marks the first epoch at which training accuracy reaches $100\%$; on a log axis this event is still easy to spot, but short post-interpolation intervals can appear visually narrow if they do not cover a large multiplicative window. To avoid duplication with the main text, we omit the path-norm panels here and focus on complementary families whose behavior further illustrates optimizer sensitivity.
The PAC–Bayes family provides a clean illustration of this theme. Under Adam, multiple bounds show slow, persistent growth after interpolation; on a log-time axis this appears as a steady positive slope across late decades of epochs, most clearly for PACBAYES_ORIG and PACBAYES_INIT. With momentum SGD, the same curves remain essentially flat within error bars once the dashed line is crossed, and on the log axis they sit nearly horizontal, emphasizing stability rather than drift.
(a) Adam
(b) SGD with momentum
Figure S10: Temporal behavior for PAC–Bayes variants on ResNet–50/FashionMNIST at fixed learning rate $0.01$. The epoch axis is logarithmic and the red dashed vertical line marks the first $100\%$ training-accuracy epoch. Adam exhibits slow post‑accuracy growth (notably in PACBAYES_ORIG and PACBAYES_INIT), whereas SGD with momentum keeps the family flat once the dashed line is crossed.
Measures tied to weight scale show the starkest divergence. In the Frobenius panel, Adam drives both the distance to initialization and the parameter norm upward almost monotonically after the model has interpolated; on a log-time axis this shows up as a persistent positive slope across late epochs. By contrast, momentum SGD leaves both traces effectively horizontal once the dashed line is passed, highlighting a stable plateau.
(a) Adam
(b) SGD with momentum
Figure S11: Frobenius distance and parameter norm through time. The epoch axis is logarithmic. Adam produces continued growth in both FRO_DIST and PARAM_NORM after $100\%$ training accuracy, while SGD with momentum holds them near a constant level.
Not every family bends under this perturbation. Both optimizers rapidly reduce the inverse‑margin surrogate early in training and then hold it near a floor. The log-time axis spreads out the initial drop, making the shared shape and timing transparent; Adam settles at a slightly higher level, but the trajectories otherwise coincide.
(a) Adam
(b) SGD with momentum
Figure S12: Inverse‑margin surrogate through time on a logarithmic epoch axis. Both optimizers shrink the measure quickly and then hover near a floor; Adam stabilizes at a slightly higher level but the overall shape is shared.
Spectral surrogates reveal a subtler but consistent imprint. Under Adam, the distance from initialization in spectral norm drifts upward over time; on log time the slope is small but positive beyond the dashed line. Under momentum SGD, the same quantity gently decreases from a plateau, appearing as a mild negative slope. The ratio FRO_OVER_SPEC also separates in level, hinting that the optimizer reshapes how mass is distributed across singular directions even when predictive performance is unchanged.
(a) Adam
(b) SGD with momentum
Figure S13: Spectral metrics through time. The epoch axis is logarithmic. Adam drives DIST_SPEC_INIT upward after interpolation, while momentum SGD yields a gentle decline; FRO_OVER_SPEC diverges slightly in level.
As a neutral reference, the VC‑dimension proxy behaves identically across optimizers by construction, and the accompanying generalization error follows the same calm trajectory. The log-time axis makes this invariance explicit: the traces remain overlapped across decades of epochs, serving as anchors that remind us not all measures are sensitive to the optimizer change.
(a) Adam
(b) SGD with momentum
Figure S14: VC‑dimension proxy and generalization error through time on a logarithmic epoch axis. The parameter‑count surrogate is identical for both optimizers and the error trajectories are similarly calm, providing a neutral reference.
Taken together, these appendix figures broaden the temporal evidence. Several measure families that depend directly on weight scale or spectrum—Frobenius norms, spectral surrogates, and parts of the PAC–Bayes suite—react strongly to an optimizer swap despite matching accuracy, while others such as margin‑based quantities and the VC proxy remain largely stable. Reading these results in aggregate helps separate measure‑intrinsic behavior from optimizer‑driven drift and, with the log-time view, clarifies whether apparent motion reflects genuine multiplicative change or merely late-stage additive updates. For temporal behavior of path norms, see Section 4, where those panels are discussed in detail.
Appendix C Proof for theorem 6.2
In this section we prove Theorem 6.2. We first reintroduce the scale invariance lemma from Arora et al. (2018b); Li & Arora (2019) which is the key source of intuition about scale-invariant networks.
Lemma C.1 (Scale-invariant networks).
If $\forall c\in\mathbb{R}^{+}$, $L(\boldsymbol{\theta})=L(c\boldsymbol{\theta})$, then
  1. 1.
    $\left\langle\nabla_{\boldsymbol{\theta}}L,\boldsymbol{\theta}\right\rangle=0$
  2. 2.
    $\left.\nabla_{\boldsymbol{\theta}}L\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}=\left.c\nabla_{\boldsymbol{\theta}}L\right|_{\boldsymbol{\theta}=c\boldsymbol{\theta}_{0}}$, for any $c>0$
Proof.
Treat $c$ as a differentiable variable. Clearly the derivative of $L$ w.r.t. $c$ is 0.
  1. 1.
    $0=\frac{\partial L}{\partial c}=\left\langle\nabla_{\boldsymbol{\theta}}L,\boldsymbol{\theta}\right\rangle$
  2. 2.
    Take gradient of the both sides of the equation $L(\boldsymbol{\theta})=L(c\boldsymbol{\theta})$ w.r.t. $\boldsymbol{\theta}$ and set $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$, we immediately arrive at the result.
We give the technical definition of the commonly used training algorithm SGD with momentum and weight decay (WD) (with respect to the L2 norm of the parameters) using a convenient form given in Li & Arora (2019), which is equivalent to the default implementation in Pytorch (Paszke et al., 2017).
Definition C.2 (SGD with momentum and WD).
At iteration $t$, given the current parameters and learning rate $(\boldsymbol{\theta}_{t-1}$, $\eta_{t-1})$, the buffered parameters and learning rate $(\boldsymbol{\theta}_{t-2}$, $\eta_{t-2})$, momentum factor $\gamma$, current loss $L(\boldsymbol{\theta}_{t-1})$ and WD factor $\lambda_{t-1}$, update the parameters $\boldsymbol{\theta}_{t}$ as the following:
$\frac{\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t-1}}{\eta_{t-1}}=\gamma\frac{\boldsymbol{\theta}_{t-1}-\boldsymbol{\theta}_{t-2}}{\eta_{t-2}}-\nabla_{\boldsymbol{\theta}}\left(L\left(\boldsymbol{\theta}_{t-1}\right)+\frac{\lambda_{t-1}}{2}\left\|\boldsymbol{\theta}_{t-1}\right\|_{2}^{2}\right)$ (21)
For the boundary conditions, it is a common practice to set $\boldsymbol{\theta}_{-1}=\boldsymbol{\theta}_{0}$ and $\eta_{-1}$ can be arbitrary.
From the above definition, it is easy to see that we can represent the state of the algorithm using a four-tuple $\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta^{\prime}},\eta^{\prime}\right)$, which stand for the parameters/learning rate at the current step and their buffers from the last step, respectively. A gradient descent step at time $t$ with momentum factor $\gamma$ and WD factor $\lambda$ can be seen as a mapping between two states:
  • A GD step with momentum and WD: $\operatorname{GD}_{t}^{\rho}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)=\left(\rho\boldsymbol{\theta}+\eta\left(\gamma\frac{\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}}{\eta^{\prime}}-\nabla L(\boldsymbol{\theta})\right),\eta,\boldsymbol{\theta},\eta\right)$
Here with WD factor being $\lambda$, $\rho$ is set to be $1-\lambda\eta$. Furthermore, we define some extra basic mappings that can be composed together to represent the temporal behavior of the algorithm.
  • Scale current parameters $\boldsymbol{\theta}$:
    $\Pi_{1}^{c}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)=\left(c\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)$;
  • Scale current LR $\eta$:
    $\Pi_{2}^{c}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)=\left(\boldsymbol{\theta},c\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)$;
  • Scale buffered parameter $\boldsymbol{\theta^{\prime}}$:
    $\Pi_{3}^{c}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)=\left(\boldsymbol{\theta},\eta,c\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)$;
  • Scale buffered LR $\eta^{\prime}$:
    $\Pi_{4}^{c}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime}\right)=\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},c\eta^{\prime}\right)$.
[…]
Definition C.3 (Equivalent states).
[…]
Here we provide an intuitive explanation of why the equivalent scaling takes the form above. If we rearrange the first term of the R.H.S. of the GD update, and assume we are operating in a regime where $\eta^{\prime}=\eta$  [^1] This just means in the eyes of the GD algorithm, the buffered LR and the current LR are the same. It does not exclude the possibility that we can still scale the current LR between GD updates. , we have
$\boldsymbol{\theta}_{\text{update}}=(\rho+\gamma)\boldsymbol{\theta}-\eta\nabla L(\boldsymbol{\theta})-\gamma\boldsymbol{\theta}^{\prime}$ (22)
In order to keep the updated parameters in the same direction, the three terms in equation 22 need to have the same scaling. From lemma C.1 we know that when $\boldsymbol{\theta}$ is scaled by $c$, the gradient $\nabla L(\boldsymbol{\theta})$ will actually be scaled by $\frac{1}{c}$. Hence for the second term $\eta\nabla L(\boldsymbol{\theta})$ to have the same amount of scaling as the first and third terms, $\eta$ has to be scaled by $c^{2}$.
The following lemma tells us that equivalent scaling commutes with GD update with momentum and WD, implying that equivalence is preserved under GD updates. Hence we are free to stack GD updates and insert equivalent scaling anywhere in a sequence of basic maps without changing the network function.
Lemma C.4 (Equivalent scaling commutes with GD).
$\forall c,\rho>0$ and $t\geq 0$,
$\text{GD}_{t}^{\rho}\circ\left[\Pi_{1}^{c}\circ\Pi_{2}^{c^{2}}\circ\Pi_{3}^{c}\circ\Pi_{4}^{c^{2}}\right]=\left[\Pi_{1}^{c}\circ\Pi_{2}^{c^{2}}\circ\Pi_{3}^{c}\circ\Pi_{4}^{c^{2}}\right]\circ\text{GD}_{t}^{\rho}.$
[…]
Proof.
For any given state $(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime})$, the L.H.S. of the equation is:
$\text{GD}_{t}^{\rho}\circ\left[\Pi_{1}^{c}\circ\Pi_{2}^{c^{2}}\circ\Pi_{3}^{c}\circ\Pi_{4}^{c^{2}}\right](\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime})$ $=\text{GD}_{t}^{\rho}(c\boldsymbol{\theta},c^{2}\eta,c\boldsymbol{\theta}^{\prime},c^{2}\eta^{\prime})$
$=\left(\rho c\boldsymbol{\theta}+c^{2}\eta\left(\gamma\frac{c\boldsymbol{\theta}-c\boldsymbol{\theta^{\prime}}}{c^{2}\eta^{\prime}}-\nabla L(c\boldsymbol{\theta})\right),c^{2}\eta,c\boldsymbol{\theta},c^{2}\eta\right)$
$=\left(c\left(\rho\boldsymbol{\theta}+\eta\left(\gamma\frac{\boldsymbol{\theta}-\boldsymbol{\theta^{\prime}}}{\eta^{\prime}}-\nabla L(\boldsymbol{\theta})\right)\right),c^{2}\eta,c\boldsymbol{\theta},c^{2}\eta\right)$
$=\left[\Pi_{1}^{c}\circ\Pi_{2}^{c^{2}}\circ\Pi_{3}^{c}\circ\Pi_{4}^{c^{2}}\right]\circ\text{GD}_{t}^{\rho}(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta^{\prime})$
Now comes the important step: in order to show the equivalence between two series of parameters with (fixed WD + fixed LR)/(exponentially decreasing WD + exponentially increasing LR), respectively, we need to rewrite $\text{GD}_{t}^{\rho}$ as a composition of itself with varying WD factor and upscaling LR, conjugated with other scaling terms that cancel with each other eventually. Here again we choose to work in the regime where the current and buffered LR are the same in the input of $\text{GD}_{t}^{\rho}$.
Lemma C.5 (Conjugated GD updates).
For any input with equal current and buffered LR $\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$ and $\forall\alpha\in\left(Z_{0},Z_{1}\right]\cup\left[Z_{2},1\right)$  [^2] Technically $\alpha$ can be larger than $1$, but in that case we will be shrinking the LR between GD steps which is not what we mainly care about here. , we have
$\text{GD}_{t}^{\rho}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)=\left[\Pi_{4}^{\alpha}\circ\Pi_{2}^{\alpha}\circ\Pi_{1}^{\alpha}\circ\text{GD}^{\beta}_{t}\circ\Pi_{2}^{\alpha^{-1}}\circ\Pi_{3}^{\alpha}\circ\Pi_{4}^{\alpha}\right]\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$
which can be written in the form of equivalent states:
$\text{GD}_{t}^{\rho}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)\stackrel{{\scriptstyle\alpha}}{{\sim}}\left[\Pi_{3}^{\alpha^{-1}}\circ\Pi_{4}^{\alpha^{-1}}\circ\Pi_{2}^{\alpha^{-1}}\circ\text{GD}_{t}^{\beta}\circ\Pi_{2}^{\alpha^{-1}}\circ\Pi_{3}^{\alpha}\circ\Pi_{4}^{\alpha}\right]\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$ (23)
where
  • $\beta=\frac{(\rho+\gamma)}{\alpha}-\frac{\gamma}{\alpha^{2}}$
  • $Z_{0}=\frac{\gamma}{1-\lambda\eta_{0}+\gamma}$
  • $Z_{1}=\frac{1+\gamma-\lambda\eta_{0}-\sqrt{(1-\gamma)^{2}-2(1+\gamma)\lambda\eta_{0}+\lambda^{2}\eta_{0}^{2}}}{2}$
  • $Z_{2}=\frac{1+\gamma-\lambda\eta_{0}+\sqrt{(1-\gamma)^{2}-2(1+\gamma)\lambda\eta_{0}+\lambda^{2}\eta_{0}^{2}}}{2}$  [^3] It’s easy to verify that $Z_{2}$ is always smaller than 1.
Proof.
We directly verify the equivalence. The R.H.S. is:
$\left[\Pi_{4}^{\alpha}\circ\Pi_{2}^{\alpha}\circ\Pi_{1}^{\alpha}\circ\text{GD}^{\beta}_{t}\circ\Pi_{2}^{\alpha^{-1}}\circ\Pi_{3}^{\alpha}\circ\Pi_{4}^{\alpha}\right]\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$
$=$ $\left[\Pi_{4}^{\alpha}\circ\Pi_{2}^{\alpha}\circ\Pi_{1}^{\alpha}\right]\circ\text{GD}^{\beta}_{t}\left(\boldsymbol{\theta},\alpha^{-1}\eta,\alpha\boldsymbol{\theta}^{\prime},\alpha\eta\right)$
$=$ $\left[\Pi_{4}^{\alpha}\circ\Pi_{2}^{\alpha}\circ\Pi_{1}^{\alpha}\right]\left(\beta\boldsymbol{\theta}+\alpha^{-1}\eta\left(\gamma\frac{\boldsymbol{\theta}-\alpha\boldsymbol{\theta}^{\prime}}{\alpha\eta}-\nabla L(\boldsymbol{\theta})\right),\alpha^{-1}\eta,\boldsymbol{\theta},\alpha^{-1}\eta\right)$
$=$ $\left(\alpha\beta\boldsymbol{\theta}+\eta\left(\gamma\frac{\boldsymbol{\theta}-\alpha\boldsymbol{\theta}^{\prime}}{\alpha\eta}-\nabla L(\boldsymbol{\theta})\right),\eta,\boldsymbol{\theta},\eta\right)$
$=$ $\left(\left(\rho+\gamma\right)\boldsymbol{\theta}-\gamma\boldsymbol{\theta}^{\prime}-\eta\nabla L(\boldsymbol{\theta}),\eta,\boldsymbol{\theta},\eta\right)$
$=$ $\left(\rho\boldsymbol{\theta}+\eta\left(\gamma\frac{\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}}{\eta}-\nabla L(\boldsymbol{\theta})\right),\eta,\boldsymbol{\theta},\eta\right)$
$=$ $\text{GD}_{t}^{\rho}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$
The range of $\alpha$ can be easily shown by combining the following two constraints and assuming remark 6.3 is true:
  • $\alpha\in\left(0,1\right];$
  • $\frac{(\rho+\gamma)}{\alpha}-\frac{\gamma}{\alpha^{2}}\in\left(0,1\right]$
Now we are ready to prove theorem 6.2.
Proof of Theorem 6.2.
From the assumption we have the following equivalence between the boundary conditions of the two series of states:
$\left(\boldsymbol{\theta}_{0},\eta_{0},\boldsymbol{\theta}_{-1},\eta_{-1}\right)$ $=\left(\boldsymbol{\theta}_{0},\eta_{0},\boldsymbol{\theta}_{0},\eta_{0}\right)$
$\left(\widetilde{\boldsymbol{\theta}}_{0},\widetilde{\eta}_{0},\widetilde{\boldsymbol{\theta}}_{-1},\widetilde{\eta}_{-1}\right)$ $=\left[\Pi_{2}^{\alpha^{-1}}\circ\Pi_{3}^{\alpha}\circ\Pi_{4}^{\alpha}\right]\left(\boldsymbol{\theta}_{0},\eta_{0},\boldsymbol{\theta}_{-1},\eta_{-1}\right)$
Lemma C.4 tells us that equivalent states are still equivalent after both being transformed by a GD step. Hence we can stack up on both sides of equation 23 for a finite number of times. i.e. for $\forall t\geq 0$, we have
$\text{GD}^{\rho}_{t-1}\circ\text{GD}^{\rho}_{t-2}\circ\cdot\circ\text{GD}^{\rho}_{0}\left(\boldsymbol{\theta}_{0},\eta_{0},\boldsymbol{\theta}_{-1},\eta_{-1}\right)$
$\stackrel{{\scriptstyle\alpha^{t}}}{{\sim}}$ $\left[\Pi_{3}^{\alpha^{-1}}\circ\Pi_{4}^{\alpha^{-1}}\circ\Pi_{2}^{\alpha^{-1}}\circ\text{GD}^{\beta}_{t-1}\circ\Pi_{2}^{\alpha^{-1}}\circ\Pi_{3}^{\alpha}\circ\Pi_{4}^{\alpha}\right]$
$\circ\cdot\circ\left[\Pi_{3}^{\alpha^{-1}}\circ\Pi_{4}^{\alpha^{-1}}\circ\Pi_{2}^{\alpha^{-1}}\right]\circ\text{GD}^{\beta}_{0}\left(\widetilde{\boldsymbol{\theta}}_{0},\widetilde{\eta}_{0},\widetilde{\boldsymbol{\theta}}_{-1},\widetilde{\eta}_{-1}\right)$
$\stackrel{{\scriptstyle\alpha^{t}}}{{\sim}}$ $\left[\Pi_{3}^{\alpha^{-1}}\circ\Pi_{4}^{\alpha^{-1}}\circ\Pi_{2}^{\alpha^{-1}}\circ\text{GD}^{\beta}_{t-1}\circ\Pi_{2}^{\alpha^{-2}}\circ\text{GD}^{\beta}_{t-2}\circ\cdot\cdot\cdot\circ\Pi_{2}^{\alpha^{-2}}\circ\text{GD}^{\beta}_{0}\right]\left(\widetilde{\boldsymbol{\theta}}_{0},\widetilde{\eta}_{0},\widetilde{\boldsymbol{\theta}}_{-1},\widetilde{\eta}_{-1}\right)$
which implies that
  • $\boldsymbol{\theta}_{t}=\alpha^{t}\widetilde{\boldsymbol{\theta}}_{t}$
  • $\widetilde{\eta}_{t}=\alpha^{-2t}\widetilde{\eta}_{0}=\alpha^{-2t-1}\eta_{0}$
  • $\widetilde{\lambda}_{t}=\frac{1-\beta}{\widetilde{\eta}_{t}}$, except for $t=0$, in which case $\widetilde{\lambda}_{0}=\frac{1-\beta+\frac{\gamma}{\alpha^{2}}-\frac{\gamma}{\alpha}}{\widetilde{\eta}_{0}}$
We note the special boundary condition for $\widetilde{\lambda}_{0}$ due to the fact that $\widetilde{\boldsymbol{\theta}}_{-1}\neq\widetilde{\boldsymbol{\theta}}_{0}$. ∎
Appendix D Additional label-corruption results: PAC–Bayes and Path Norms
This appendix gathers the label–corruption results referenced in the main text for both PAC–Bayes–style measures and Path Norms.
PAC–Bayes measures.
Under Adam, lowering the learning rate from $10^{-2}$ to $10^{-3}$ transforms a clean, steadily rising ramp (e.g., PACBAYES_ORIG from $\sim 3.6$ past $4.7$) into a higher–lying but shallower U–shape centred around $\sim 8$. At fixed learning rate $10^{-2}$, swapping Adam for SGD produces a striking level shift: SGD yields a high plateau ($\sim 12$$13$) with little curvature, while Adam traces a low, clearly increasing arc ($\sim 3.6$$4.7$). Even within the PAC–Bayes family, members respond differently by optimizer: PACBAYES_INIT grows aggressively under Adam but only mildly under SGD. These patterns illustrate an (often) insensitive or contradictory relationship between the measure and increasing label corruption—another facet of fragility.
(a) Adam, LR =0.01=0.01
(b) Adam, LR =0.001=0.001
Figure S15: PAC–Bayes measures vs. label corruption under Adam. With $\mathrm{LR}{=}10^{-2}$, the family rises steadily with corruption; with $\mathrm{LR}{=}10^{-3}$ it sits higher overall and shows a shallow U–shape. All panels use 10,000 training samples.
(a) Adam, LR =0.01=0.01
(b) SGD (momentum), LR =0.01=0.01
Figure S16: PAC–Bayes measures vs. label corruption at fixed LR ($10^{-2}$), swapping only the optimizer. Adam yields a low, steadily rising family (e.g., PACBAYES_ORIG $\approx 3.6$$4.7$), whereas SGD holds a high plateau ($\sim 12$$13$) with minimal curvature; PACBAYES_INIT grows far more under Adam than under SGD. All panels use 10,000 training samples.
Path norms under SGD (momentum).
Mirroring the Adam case in the main text, SGD with momentum exhibits an equally striking flip (Fig. S17): at $\mathrm{LR}{=}10^{-3}$, the path-norm trajectory lives in the $10^{4}$$10^{5}$ band and decays steadily with corruption; at $\mathrm{LR}{=}10^{-1}$, the scale crashes to nearly zero and the curve climbs monotonically. Trend and scale both invert. If a reader tried to infer “harder data $\Rightarrow$ larger path norm” from one panel, the other would immediately contradict it.
(a) LR =0.1=0.1
(b) LR =0.001=0.001
Figure S17: Path norms vs. label corruption with SGD (momentum); panels differ only in learning rate. At $\mathrm{LR}{=}10^{-3}$ the curve sits in $10^{4}$$10^{5}$ and decays with corruption; at $\mathrm{LR}{=}10^{-1}$ it lives near zero and rises. Both the direction and the dynamic range flip—another instance of qualitative mismatch across a minimal training change. All panels use 10,000 training samples.
Appendix E Post-mortem vs. ML-PACBayes
We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of Valle-Pérez & Louis (2020). The bound controls the test error of a hypothesis sampled from a Bayesian posterior over functions (not parameters); the key capacity term is the marginal likelihood (Bayesian evidence) of the data under the architecture’s Gaussian‑process (GP) limit. All figure numbers below refer to Valle-Pérez & Louis (2020).
Definition E.1 (Marginal-likelihood PAC–Bayes bound).
Consider binary classification with data distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$ and a hypothesis space $\mathcal{H}$ of functions $h:\mathcal{X}\to\{0,1\}$. Let $S\sim\mathcal{D}^{n}$ be a training set of size $n\geq 2$, and let $P$ be a prior on $\mathcal{H}$. Define the consistency set
$C(S)\;:=\;\{\,h\in\mathcal{H}\;:\;\hat{\varepsilon}(h,S)=0\,\},$
where $\hat{\varepsilon}(h,S)=\tfrac{1}{n}\sum_{(x,y)\in S}\mathbf{1}\{h(x)\neq y\}$ is the empirical 0–1 error. The (realizable) Bayesian posterior supported on $C(S)$ is
$Q(h)\;=\;\begin{cases}\dfrac{P(h)}{P(C(S))},&h\in C(S),\\[4.0pt] 0,&\text{otherwise},\end{cases}$
where $P(C(S))=\sum_{h\in C(S)}P(h)$ is the marginal likelihood (Bayesian evidence) of $S$.
For any confidence levels $\delta,\gamma\in(0,1]$, with probability at least $1-\delta$ over $S\sim\mathcal{D}^{n}$ and, conditional on $S$, with probability at least $1-\gamma$ over $h\sim Q$, the generalization error $\varepsilon(h)=\Pr_{(x,y)\sim\mathcal{D}}[h(x)\neq y]$ satisfies
$\ln\!\bigl(1-\varepsilon(h)\bigr)\;<\;\frac{\ln\!\tfrac{1}{P(C(S))}+\ln n+\ln\!\tfrac{1}{\delta}+\ln\!\tfrac{1}{\gamma}}{\,n-1\,}.$ (24)
Remark E.2 (binary vs. multiclass).
We state the bound for binary classification for clarity; the function‑space marginal‑likelihood perspective extends to multiclass via standard reductions (e.g., one‑vs‑rest) or multiclass likelihoods.
What our fragility checks reveal—and what the ML-PACBayes bound gets right.
We evaluated each measure against consequential perturbations: dataset difficulty, data-size scaling, training‑pipeline changes, and temporal behavior (overtraining). Post‑mortem measures (sharpness/flatness, norm/margin proxies, compression‑style, and deterministic PAC–Bayes surrogates) often fail at least one stressor; in our runs, even strong performers such as PACBAYES_orig failed key tests. By contrast, the GP‑based bound tracks the function‑level regularities we care about:
  • Data complexity. The marginal‑likelihood bound tracks dataset difficulty: it increases with label corruption and preserves the canonical ordering MNIST << Fashion‑MNIST << CIFAR‑10; see Figure 1 in (Valle-Pérez & Louis, 2020). In our audits, several post‑mortem measures—including PACBAYES_orig—either flattened under corruption or were not able to catch the ordering, even when the test error did.
  • Learning‑curve scaling. The marginal‑likelihood bound tracks learning‑curve scaling: it mirrors the empirical power‑law in $n$ and clusters exponents primarily by dataset; see Figures 2–5 in (Valle-Pérez & Louis, 2020). By contrast, some post‑mortem measures barely moved with $n$; others bent the wrong way or mixed optimizer effects with data effects (cf. (Nagarajan & Kolter, 2019)).
  • Temporal behavior (overtraining invariance). The marginal‑likelihood bound tracks temporal invariance: once $S$ is fixed and training has interpolated ($\hat{\varepsilon}=0$), the bound depends only on $P(C(S))$ and nn—it is agnostic to how long or by which path the parameters were trained. In our temporal‑behavior experiments, several norm‑based post‑mortem measures (e.g., $\ell_{2}$/spectral/path norms) kept growing during overtraining while the generalization error stayed roughly unchanged; the GP bound remained stable. (This echoes broader observations that longer training can leave test error flat or improved (Hoffer et al., 2017), and that norms can diverge under separable losses without hurting classification error (Soudry et al., 2018).)
  • Training‑pipeline changes. Many post‑mortem measures swing with optimizer, batch size, explicit/implicit regularization, and early stopping—sometimes tracking curvature artifacts rather than out‑of‑sample error (e.g., Keskar et al., 2017; Hochreiter & Schmidhuber, 1997; Jiang et al., 2019b; Dziugaite et al., 2020b). The GP marginal‑likelihood bound is, by construction, insensitive to these knobs: it depends on the architecture’s function prior and the data, not on the path taken through parameter space. We view this invariance as a virtue for a predictor; the bound is not designed to explain differences caused by optimizer choice and should not be judged on that criterion.
These properties are consistent with our optimizer‑swap and pixel‑permutation stress tests (Apps. B and F) and with our scale‑symmetry Exp++ dial, which only alters parameter scale (App. G).
Why does the function‑space route outperform post‑mortems?
Two ingredients stand out. First, the bound evaluates an Occam factor in function space: architectures that place high prior mass on data‑compatible functions earn better evidence, naturally capturing dataset ordering and learning‑curve slopes. Second, there is a credible external yardstick: infinite‑width GP predictors (NNGP/NTK) often quantitatively predict finite‑width DNN generalization trends (Lee et al., 2020a). Together, these explain why Figures 1–5 show greater stability for the GP‑based approach than we observe for post‑hoc, parameter‑space measures.
So why do post‑mortems underperform?
The live hypothesis is that they measure the wrong object. Post‑mortem scores probe properties of a single trained parameter vector (curvature, norms, compressibility), entangling optimization details and reparameterization choices with generalization. That enterprise is important—and the community has argued forcefully that such post‑training diagnostics deserve attention (e.g., compression and robustness perspectives in Arora et al., 2018a, 2019; Jiang et al., 2019b; Dziugaite et al., 2020b)—but our evidence suggests their predictions are fragile under routine perturbations. By contrast, the GP bound targets the distribution over functions implied by architecture and data.
Open question.
Even if the GP‑based predictor/bound wins these stress tests, finite‑width networks can outperform their GP limits; yet the GP still predicts a striking fraction of performance trends (Figures 2–5). Closing this gap—by designing reparameterization‑invariant, data‑aware post‑mortem diagnostics that inherit the Occam flavour of marginal likelihood—remains open.
Appendix F Stress-testing generalization measures with pixel permutations
A useful generalization measure should pass two basic stress tests: (i) it should be insensitive to symmetry-preserving transformations that do not change the intrinsic task, yet (ii) sensitive when task-relevant information is destroyed. Pixel permutations provide such a testbed: applying the same permutation to the training and test sets preserves label–input relationships (a symmetry for fully connected networks), whereas applying independent random permutations to train and test destroys spatial structure and any usable signal.
We evaluate three families of measures alongside test error on MNIST with two architectures: a fully connected network (FCN) and a ResNet-50. We vary the optimizer (sgd vs. adam) and the early-stopping criterion (best cross-entropy “CE” vs. best accuracy “Acc”) to expose hyperparameter sensitivity. Results are summarized in Tables 1 and 2.
FCN (permutation symmetry holds).
When the same pixel permutation is applied to both train and test, the FCN effectively sees an unchanged task. As expected, test error remains essentially constant ($\approx 0.03$ across optimizers/stopping), and the PAC-Bayesian marginal likelihood (ML) bound is flat ($0.142$ throughout). This indicates desirable robustness to symmetry-preserving changes. In contrast, under independent random permutations, test error rises to $\approx 0.49$$0.50$ (near random guessing), and the ML bound increases to 0.5320.532—appropriately reflecting reduced learnability. The path norm and the original PAC-Bayes variant (PACBAYES-ORIG) trend upward in this harder setting, but they also exhibit large spread across optimizers/stopping; e.g., the path norm spans from $0.105$ to $6.632$ under random shuffling, smaller than some unshuffled cases, revealing fragility to seemingly minor training choices.
ResNet-50 (permutation symmetry broken).
Because convolutional inductive biases depend on spatial locality, even a single shared permutation distorts the data geometry and degrades performance (test error increases from $\approx 0.014$$0.025$ to $\approx 0.041$$0.060$; the ML bound rises from $0.101$ to $0.170$). With independent random permutations, test error again moves to $\approx 0.44$$0.46$ and the ML bound to $\approx 0.504$. The path norm is particularly volatile here, spanning orders of magnitude across training choices (e.g., $0.016$ vs. $3375.355$), underscoring substantial measure fragility.
Takeaways.
Under symmetry-preserving changes (FCN + same-permutation), a robust measure should not move; under signal destruction (independent permutations), it should reflect the loss of learnability. The PAC-Bayesian ML bound behaves in this manner in both architectures, whereas the path norm and PACBAYES-ORIG can vary dramatically with optimizer/stopping, masking the underlying data effect. Reporting such sensitivity is crucial when proposing or comparing generalization measures, in line with our paper’s emphasis on diagnosing and documenting fragility. [^4] While label corruption is another standard knob for data complexity, here we focus on pixel permutations to isolate architectural symmetry vs. information destruction.
Table 1: Pixel permutations with MNIST + FCN (training set size n=10,000n{=}10{,}000). Same-permutation preserves the FCN’s permutation symmetry; independent (random) permutations destroy learnable signal.
Optimizer / stopping Test err. PAC-Bayes ML Path norm PACBAYES-ORIG
Unshuffled pixels
SGD / CE $0.033$ $0.142$ $0.080$ $1.534$
ADAM / CE $0.034$ $0.142$ $0.173$ $1.208$
SGD / Acc $0.032$ $0.142$ $0.081$ $1.535$
ADAM / Acc $0.031$ $0.142$ $0.379$ $1.187$
Same permutation on train & test
SGD / CE $0.032$ $0.142$ $0.080$ $1.558$
ADAM / CE $0.034$ $0.142$ $0.184$ $1.202$
SGD / Acc $0.032$ $0.142$ $0.081$ $1.534$
ADAM / Acc $0.032$ $0.142$ $0.314$ $1.183$
Independent random permutations (train & test)
SGD / CE $0.497$ $0.532$ $0.124$ $2.232$
ADAM / CE $0.492$ $0.532$ $5.736$ $1.791$
SGD / Acc $0.498$ $0.532$ $0.105$ $2.400$
ADAM / Acc $0.488$ $0.532$ $6.632$ $1.837$
  • CE: best checkpoint by cross-entropy; Acc: best checkpoint by accuracy.
Table 2: Pixel permutations with MNIST + ResNet-50 (n=10,000n{=}10{,}000). Convolutional inductive biases depend on spatial locality, so even a single shared permutation harms performance.
Optimizer / stopping Test err. PAC-Bayes ML Path norm PACBAYES-ORIG
Unshuffled pixels
SGD / CE $0.022$ $0.101$ $2980.415$ $15.396$
ADAM / CE $0.016$ $0.101$ $0.173$ $4.498$
SGD / Acc $0.025$ $0.101$ $2972.887$ $15.125$
ADAM / Acc $0.014$ $0.101$ $0.022$ $4.417$
Same permutation on train & test
SGD / CE $0.060$ $0.170$ $3375.355$ $13.534$
ADAM / CE $0.046$ $0.170$ $0.016$ $4.074$
SGD / Acc $0.056$ $0.170$ $3001.528$ $12.529$
ADAM / Acc $0.041$ $0.170$ $0.021$ $3.909$
Independent random permutations (train & test)
SGD / CE $0.460$ $0.504$ $116.602$ $13.77$
ADAM / CE $0.437$ $0.504$ $0.081$ $4.157$
SGD / Acc $0.460$ $0.504$ $67.497$ $12.475$
ADAM / Acc $0.437$ $0.504$ $0.084$ $4.119$
  • CE: best checkpoint by cross-entropy; Acc: best checkpoint by accuracy.
Appendix G Exp++ in scale-invariant nets: protocol and results
We instantiate the symmetry in a fully connected, scale‑invariant network (normalization after each hidden linear layer; the final linear readout is frozen), trained on MNIST with SGD + momentum. We sweep a single control, the Exp++ multiplicative factor $\alpha$ from Theorem 6.2, which makes the learning rate grow exponentially across steps, and repeat the sweep both without and with weight decay. For each run we log parameter‑space PAC–Bayes proxies (including magnitude‑aware variants), path‑norm proxies (with/without margin normalization), and test error. The stopping rule is cross‑entropy unless stated otherwise.
Findings. Across both regimes (no WD and with WD) the test‑error trace barely moves as $\alpha$ grows, yet the magnitude‑sensitive diagnostics explode on a log scale. In particular, the PAC–Bayes family and the path‑norm proxies rise by many orders of magnitude even though the error curve hugs a dashed 10% reference. This is exactly what the equivalence predicts in a scale‑invariant net: Exp++ changes parameter scale, not the predictor—exposing the fragility of magnitude‑dependent measures (notably PACBAYES_orig) in this setting.
Figure S18: Exp++ in a scale‑invariant FCN (no WD; CE stopping). As the Exp++ factor increases, PAC–Bayes proxies (left) and path‑norm proxies (right) swell by orders of magnitude, while the test‑error curve (right axis; dashed 10% line) remains essentially flat.
Figure S19: Exp++ with weight decay (CE stopping). The qualitative picture persists with WD: magnitude‑sensitive PAC–Bayes and path‑norm diagnostics climb sharply with the Exp++ factor, but the predictor’s test error barely changes.
Reading the panels. In both Figure S18 (no WD) and Figure S19 (with WD), left panels aggregate PAC–Bayes proxies; right panels show path‑norm proxies. Axes are logarithmic for the measures and linear for the test error (right axis); the dashed horizontal line marks 10% test error for visual reference. Together, the panels illustrate a clean separation between parameter scale (which Exp++ manipulates) and predictive behavior (which stays essentially fixed).

### v2

@@
 Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends.
-In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behaviour.
-For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalisation measures like the path norm.
+In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behavior.
+For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalization measures like the path norm.
 We also identify subtler forms of fragility.
@@
 However, it completely fails to capture differences in data complexity across learning curves.
-This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behaviour, in learning curves, but which is not a post-mortem measure.
-Beyond demonstrating that many bounds - such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates- are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
+This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behavior, in learning curves, but which is not a post-mortem measure.
+Beyond demonstrating that many bounds—such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC-Bayes surrogates—are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
 Machine Learning, ICML
@@
 For deep networks the same aspiration has driven a profusion of bounds.
-While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behaviour of deep neural networks (DNNs).
+While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs).
 These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds.
-Large-scale comparisons have compared the behaviour of these measures, albeit often only at the sign level, e.g.
+Large-scale comparisons have compared the behavior of these measures, albeit often only at the sign level, e.g.
 does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
@@
 This suggests that generalization measures should carefully track key invariances present in the underlying DNNs.
-The widespread fragility we observe suggests that some generalisation measures may be showing correlation in some limited settings that may not be related to causation.
+The widespread fragility we observe suggests that some generalization measures may be showing correlation in some limited settings that may not be related to causation.
 Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
@@
 Both drop faster than the underlying ResNet‑50 test error.
-Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behaviour.
+Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior.
 A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows.
-These widely different behaviours arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
-The contrasting behaviour of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression.
+These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
+The contrasting behavior of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression.
 In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer.
@@
 This illustrates the kind of fragility our framework urges authors to investigate and report when treating bounds as measures.
-Finally, the last panel in figure 3 illustrates how the function‑space ML‑PACBayes measure provides a non-vacuous bound that tracks data-complexity (see also Appendix E).
+Finally, the last panel in figure 3 illustrates how the function‑space ML‑PACBayes measure provides a non-vacuous bound that tracks data-complexity (see also Appendix F).
 This performance is not mirrored by several PAC–Bayes parameter‑space surrogates, which on the one hand are much less sensitive to training hyperparameters than the path norm measures are, but on the other hand exhibit a different more subtle kind of fragility; they are typically insensitive to label corruption in standard setups (see Appendix D).
-We call this behaviour a fragility because it reflects a serious failing of these measures.
+We call this behavior a fragility because it reflects a serious failing of these measures.
 Figure 3: Label‑corruption sweeps with post‑mortem norms vs.
@@
 By contrast, the third panel (ML‑PACBayes) is both tight (close to the corresponding error curves) and data‑aware: it preserves the MNIST << FashionMNIST << CIFAR‑10 ordering and reflects the different slopes.
-See Appendix E for the bound’s definition and a fuller discussion of why this function‑space quantity captures dataset difficulty while post-mortem parameter‑space PAC-Bayes surrogates often do not.
+See Appendix F for the bound’s definition and a fuller discussion of why this function‑space quantity captures dataset difficulty while post-mortem parameter‑space PAC-Bayes surrogates often do not.
 There has been substantial recent interest in the study of scaling laws for deep neural networks (DNNs).
@@
 In App.
-F, we move beyond label corruption and dataset swaps, and also probe symmetry‑preserving vs.
+G, we move beyond label corruption and dataset swaps, and also probe symmetry‑preserving vs.
 signal‑destroying data transforms via pixel permutations.
@@
 Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures.
-Empirical details and results are deferred to Appendix G.
+Empirical details and results are deferred to Appendix H.
 Definition 6.1 (Scale invariant neural networks).
@@
 In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures.
-Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix G.
+Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix H.
 7 Quantitative fragility: a scale‑free instability score
@@
 Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
+Full experimental details, formal algorithms and tabulated results can be found in Appendix E.
 8 Conclusion and discussion
@@
 Optimizer sensitivity as a feature.
-It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalisation measures.
+It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalization measures.
 This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019).
@@
 Limits of function–space GP references.
-A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix E) overstates its practical authority.
+A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix F) overstates its practical authority.
 Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination.
@@
 All panels use 10,000 training samples.
-Appendix E Post-mortem vs. ML-PACBayes
+Appendix E Quantitative fragility: detailed results
+This appendix extends Section 7.
+We first summarize the experimental setup, and then give concise, implementable procedures for our two instability scores: CMS and its seed‑adjusted variant eCMS.
+We close with a brief reading guide that highlights what can be learned from the summary tables at Table 1 ($\delta=0.01$), Table 2 ($\delta=0.02$), and Table 3 ($\delta=0.05$).
+For full definitions of the measures listed in those tables, see (Dziugaite et al., 2020b).
+Experimental setup.
+We study two datasets—FashionMNIST and CIFAR‑10—across three architectures: Network‑in‑Network (NiN) (Lin et al., 2013), ResNet‑50 (RN50), and DenseNet‑121 (DN121).
+The training pipeline varies along three axes: (i) learning rate $\texttt{lr}\in\{0.001,0.0032,0.0063,0.01,0.0158,0.05,0.1\}$; (ii) optimizer $\in\{\textsc{Adam},\textsc{SGD}+\textsc{momentum}\}$; and (iii) stopping criterion: either the first epoch reaching $100\%$ training accuracy, or the first epoch whose training cross‑entropy drops below $0.01$.
+Each hyperparameter configuration is replicated with $8$ random seeds.
+We compute scores within each (dataset, architecture) group at absolute error tolerances $\delta\in\{0.01,0.02,0.05\}$, using natural logs and reporting medians.
+Using the procedures in practice.
+Algorithms 1 and 2 specify the exact pair‑construction and aggregation steps used throughout our study.
+In all tables, we apply these procedures within each (dataset, architecture) group and then report group‑level medians; if a required pair set is empty for a group, the statistic is marked Undefined and the group is omitted from the across‑group aggregation.
+When the number of eligible pairs is large, we uniformly subsample up to a fixed budget $K$ before taking the median, following the same steps as in Algorithms 12.
+Algorithm 1 CMS (Conditional Measure Spread)
+1:Runs $\{(C_{r},\widehat{\varepsilon}_{r})\}_{r=1}^{n}$ with $C_{r}>0$; tolerance $\delta>0$; optional pair budget $K$.
+[…]
+3:Sort indices so that $\widehat{\varepsilon}_{(1)}\leq\cdots\leq\widehat{\varepsilon}_{(n)}$.
+[…]
+5:for $i\leftarrow 1$ to $n$ do
+6:  $j\leftarrow i+1$
+7:  while $j\leq n$ and $\widehat{\varepsilon}_{(j)}-\widehat{\varepsilon}_{(i)}\leq\delta$ do
+8:    append $\bigl|\log C_{(i)}-\log C_{(j)}\bigr|$ to $L$
+9:    $j\leftarrow j+1$
+10:if $|L|=0$ then return Undefined {no close‑error pairs}
+11:if $|L|>K$ then
+12:  uniformly subsample $K$ elements of $L$
+13:return $\operatorname{median}(L)$
+Algorithm 2 eCMS (seed‑adjusted CMS)
+1:Runs $\{(C_{r},\widehat{\varepsilon}_{r},H_{r},\textsf{seed}_{r})\}_{r=1}^{n}$ with $C_{r}>0$; tolerance $\delta>0$; optional pair budget $K$.
+[…]
+3:Sort indices so that $\widehat{\varepsilon}_{(1)}\leq\cdots\leq\widehat{\varepsilon}_{(n)}$.
+[…]
+5:for $i\leftarrow 1$ to $n$ do
+6:  $j\leftarrow i+1$
+7:  while $j\leq n$ and $\widehat{\varepsilon}_{(j)}-\widehat{\varepsilon}_{(i)}\leq\delta$ do
+8:    if $H_{(j)}=H_{(i)}$ and $\textsf{seed}_{(j)}\neq\textsf{seed}_{(i)}$ then
+9:     append $\bigl|\log C_{(i)}-\log C_{(j)}\bigr|$ to $L_{\text{seed}}$
+10:    else if $H_{(j)}\neq H_{(i)}$ then
+11:     append $\bigl|\log C_{(i)}-\log C_{(j)}\bigr|$ to $L_{\text{inter}}$
+12:    $j\leftarrow j+1$
+13:if $|L_{\text{inter}}|=0$ or $|L_{\text{seed}}|=0$ then return Undefined
+14:if $|L_{\text{seed}}|>K$ then
+15:  uniformly subsample $K$ elements of $L_{\text{seed}}$
+16:if $|L_{\text{inter}}|>K$ then
+17:  uniformly subsample $K$ elements of $L_{\text{inter}}$
+[…]
+[…]
+20:return $\max\{0,\ \mathrm{CMS}^{\text{inter}}_{\delta}-\mathrm{CMS}^{\text{seed}}_{\delta}\}$
+What the tables show.
+The summaries in Table 1, Table 2, and Table 3 reveal clear, consistent patterns:
+  • Stable control.
+The parameter‑count proxy (PARAMS) is uniformly flat (CMS =0=0, eCMS =0=0) across datasets, architectures, and tolerances; see the first row of Table 1 and its counterparts at $\delta=0.02$ and $\delta=0.05$.
+  • PAC–Bayes surrogates split.
+The classical parameter‑space PACBAYES_ORIG exhibits substantial instability (elevated CMS and eCMS), often increasing as the tolerance widens (compare the “ORIG” rows across Tables 1 and 2).
+Magnitude‑aware variants (e.g., PACBAYES_MAG_*) are consistently lower, though they remain sensitive on FashionMNIST with deeper architectures.
+  • Norm families are fragile.
+Parameter and path norms already show sizeable spreads at $\delta=0.01$ and inflate further at $\delta=0.02$$0.05$.
+Frobenius and spectral distances from initialization are larger still and consistently among the most unstable rows across tables.
+  • Product/aggregate spectral–Frobenius measures are outliers.
+Aggregate/product metrics (e.g., LOG_PROD_OF_SPEC, LOG_SPEC_* and Frobenius analogues) exhibit spreads that are orders of magnitude larger than other families at $\delta=0.01$, and this gap persists at $\delta=0.02$ and $\delta=0.05$.
+  • Beyond seeds.
+For many families (e.g., PACBAYES_ORIG, PARAM_NORM, FRO_DIST) we observe $\mathrm{eCMS}>\mathrm{CMS}$ across tolerances, indicating that a substantial portion of the jumpiness is attributable to hyperparameter changes rather than random seeds; exceptions exist for specific (dataset, architecture) cells.
+  • Where fragility concentrates.
+Per‑group columns indicate that the largest spreads tend to cluster on FashionMNIST with RN50/DN121, while CIFAR‑10/NiN often shows the smallest values for the same measure (compare group subcolumns across tables).
+Table 1: CMS and eCMS summary $\delta=0.01$
+Measure $\mathrm{CMS}^{\mathrm{med}}_{\delta}/\mathrm{eCMS}^{\mathrm{med}}_{\delta}$ C10/DN121 C10/NiN C10/RN50 FMN/DN121 FMN/NiN FMN/RN50
+PARAMS $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
+PACBAYES_MAG_ORIG $0.084$ $0.088$ $0.092$ $0.136$ $0.029$ $0.031$ $0.060$ $0.064$ $0.109$ $0.119$ $0.076$ $0.075$ $0.097$ $0.102$
+PACBAYES_MAG_INIT $0.103$ $0.114$ $0.129$ $0.179$ $0.038$ $0.041$ $0.086$ $0.102$ $0.120$ $0.126$ $0.065$ $0.062$ $0.164$ $0.159$
+FRO_OVER_SPEC $0.135$ $0.171$ $0.085$ $0.125$ $0.071$ $0.077$ $0.095$ $0.169$ $0.239$ $0.265$ $0.174$ $0.173$ $0.323$ $0.352$
+INVERSE_MARGIN $0.207$ $0.194$ $0.202$ $0.201$ $0.111$ $0.115$ $0.213$ $0.186$ $0.218$ $0.204$ $0.149$ $0.133$ $0.240$ $0.218$
+PACBAYES_MAG_FLATNESS $0.238$ $0.284$ $0.442$ $0.627$ $0.155$ $0.157$ $0.234$ $0.276$ $0.242$ $0.291$ $0.184$ $0.159$ $0.436$ $0.444$
+PACBAYES_FLATNESS $0.419$ $0.503$ $0.384$ $0.534$ $0.223$ $0.259$ $0.337$ $0.415$ $0.455$ $0.486$ $0.606$ $0.605$ $0.519$ $0.519$
+PACBAYES_ORIG $0.516$ $0.638$ $0.514$ $0.769$ $0.412$ $0.591$ $0.518$ $0.658$ $0.438$ $0.470$ $0.808$ $0.815$ $0.582$ $0.618$
+SUM_OF_FRO_OVER_MARGIN $0.647$ $0.903$ $0.596$ $0.899$ $0.470$ $0.555$ $0.699$ $0.907$ $0.539$ $0.615$ $0.863$ $1.081$ $0.889$ $1.001$
+SUM_OF_FRO $0.650$ $0.903$ $0.597$ $0.898$ $0.480$ $0.564$ $0.703$ $0.909$ $0.540$ $0.615$ $0.876$ $1.102$ $0.888$ $0.999$
+PATH_NORM_OVER_MARGIN $0.945$ $1.014$ $0.979$ $1.108$ $0.192$ $0.153$ $0.910$ $0.919$ $1.453$ $1.474$ $0.416$ $0.379$ $1.557$ $1.484$
+SUM_OF_SPEC_OVER_MARGIN $0.856$ $1.062$ $0.667$ $0.984$ $0.511$ $0.605$ $0.926$ $1.140$ $0.785$ $0.869$ $1.100$ $1.202$ $1.327$ $1.404$
+SUM_OF_SPEC $0.857$ $1.064$ $0.669$ $0.985$ $0.518$ $0.613$ $0.928$ $1.142$ $0.786$ $0.870$ $1.124$ $1.221$ $1.329$ $1.404$
+PACBAYES_INIT $0.965$ $1.066$ $0.718$ $0.944$ $0.427$ $0.621$ $0.763$ $0.862$ $1.231$ $1.331$ $1.168$ $1.187$ $1.189$ $1.224$
+PARAM_NORM $0.895$ $1.102$ $0.956$ $1.356$ $0.601$ $0.735$ $0.798$ $1.013$ $0.834$ $0.977$ $1.239$ $1.322$ $1.047$ $1.192$
+PATH_NORM $1.073$ $1.184$ $1.147$ $1.320$ $0.266$ $0.252$ $0.998$ $1.047$ $1.540$ $1.538$ $0.494$ $0.457$ $1.560$ $1.545$
+DIST_SPEC_INIT $1.371$ $1.567$ $1.084$ $1.467$ $0.612$ $0.737$ $1.151$ $1.378$ $1.840$ $1.997$ $1.590$ $1.668$ $1.748$ $1.848$
+FRO_DIST $1.405$ $1.612$ $1.071$ $1.471$ $0.633$ $0.757$ $1.106$ $1.282$ $1.718$ $1.838$ $1.705$ $1.753$ $1.744$ $1.835$
+PROD_OF_FRO $46.133$ $55.352$ $144.464$ $217.241$ $3.360$ $3.948$ $40.746$ $52.739$ $130.638$ $148.734$ $6.131$ $7.716$ $51.520$ $57.964$
+PROD_OF_FRO_OVER_MARGIN $46.051$ $55.358$ $144.118$ $217.564$ $3.289$ $3.886$ $40.567$ $52.634$ $130.509$ $148.827$ $6.042$ $7.566$ $51.535$ $58.082$
+SPEC_ORIG_MAIN $65.140$ $73.373$ $161.340$ $237.870$ $3.553$ $4.191$ $53.616$ $65.936$ $189.721$ $209.791$ $7.410$ $8.277$ $76.664$ $80.810$
+PROD_OF_SPEC_OVER_MARGIN $65.339$ $73.798$ $161.349$ $238.066$ $3.578$ $4.234$ $53.711$ $66.135$ $190.061$ $210.406$ $7.698$ $8.414$ $76.967$ $81.461$
+SPEC_INIT_MAIN $65.299$ $73.805$ $161.472$ $238.038$ $3.646$ $4.218$ $53.702$ $66.062$ $190.120$ $210.831$ $7.920$ $8.422$ $76.897$ $81.548$
+PROD_OF_SPEC $65.467$ $73.835$ $161.834$ $238.429$ $3.628$ $4.291$ $53.848$ $66.234$ $190.328$ $210.514$ $7.868$ $8.546$ $77.087$ $81.435$
+Top numbers: CMS; bottom numbers: eCMS.
+Table 2: CMS and eCMS summary $\delta=0.02$
+Measure $\mathrm{CMS}^{\mathrm{med}}_{\delta}/\mathrm{eCMS}^{\mathrm{med}}_{\delta}$ C10/DN121 C10/NiN C10/RN50 FMN/DN121 FMN/NiN FMN/RN50
+PARAMS $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
+PACBAYES_MAG_ORIG $0.094$ $0.095$ $0.106$ $0.137$ $0.046$ $0.047$ $0.064$ $0.068$ $0.107$ $0.114$ $0.082$ $0.079$ $0.110$ $0.111$
+PACBAYES_MAG_INIT $0.118$ $0.139$ $0.139$ $0.170$ $0.049$ $0.047$ $0.097$ $0.108$ $0.173$ $0.174$ $0.072$ $0.068$ $0.199$ $0.187$
+FRO_OVER_SPEC $0.165$ $0.183$ $0.100$ $0.126$ $0.088$ $0.090$ $0.092$ $0.135$ $0.270$ $0.292$ $0.230$ $0.232$ $0.410$ $0.419$
+INVERSE_MARGIN $0.227$ $0.204$ $0.224$ $0.205$ $0.122$ $0.114$ $0.247$ $0.219$ $0.230$ $0.203$ $0.147$ $0.128$ $0.251$ $0.218$
+PACBAYES_MAG_FLATNESS $0.244$ $0.253$ $0.470$ $0.597$ $0.190$ $0.188$ $0.261$ $0.283$ $0.228$ $0.224$ $0.170$ $0.143$ $0.413$ $0.401$
+PACBAYES_FLATNESS $0.486$ $0.535$ $0.431$ $0.516$ $0.328$ $0.337$ $0.352$ $0.398$ $0.542$ $0.553$ $0.746$ $0.731$ $0.607$ $0.603$
+PACBAYES_ORIG $0.557$ $0.697$ $0.581$ $0.761$ $0.768$ $0.797$ $0.533$ $0.644$ $0.438$ $0.437$ $0.755$ $0.749$ $0.503$ $0.490$
+SUM_OF_FRO $0.695$ $0.872$ $0.657$ $0.868$ $0.648$ $0.797$ $0.733$ $0.875$ $0.534$ $0.586$ $0.930$ $1.097$ $0.842$ $0.921$
+SUM_OF_FRO_OVER_MARGIN $0.694$ $0.872$ $0.656$ $0.869$ $0.646$ $0.791$ $0.733$ $0.875$ $0.533$ $0.586$ $0.913$ $1.076$ $0.839$ $0.923$
+PATH_NORM_OVER_MARGIN $1.023$ $1.008$ $1.037$ $1.043$ $0.252$ $0.204$ $1.009$ $0.973$ $1.928$ $1.891$ $0.582$ $0.546$ $2.083$ $2.017$
+SUM_OF_SPEC_OVER_MARGIN $0.882$ $1.040$ $0.735$ $0.961$ $0.772$ $0.891$ $0.952$ $1.120$ $0.811$ $0.870$ $1.162$ $1.221$ $1.365$ $1.389$
+SUM_OF_SPEC $0.884$ $1.041$ $0.734$ $0.961$ $0.774$ $0.900$ $0.956$ $1.120$ $0.813$ $0.870$ $1.182$ $1.240$ $1.364$ $1.387$
+PARAM_NORM $0.960$ $1.064$ $1.033$ $1.318$ $0.935$ $1.014$ $0.822$ $0.988$ $0.781$ $0.928$ $1.282$ $1.323$ $0.985$ $1.114$
+PACBAYES_INIT $1.052$ $1.105$ $0.818$ $0.947$ $0.850$ $0.959$ $0.777$ $0.846$ $1.438$ $1.504$ $1.254$ $1.250$ $1.390$ $1.394$
+PATH_NORM $1.158$ $1.182$ $1.208$ $1.268$ $0.349$ $0.323$ $1.107$ $1.096$ $1.914$ $1.874$ $0.623$ $0.580$ $2.116$ $2.086$
+DIST_SPEC_INIT $1.482$ $1.629$ $1.227$ $1.452$ $0.998$ $1.117$ $1.190$ $1.383$ $2.132$ $2.225$ $1.738$ $1.806$ $2.042$ $2.061$
+FRO_DIST $1.537$ $1.666$ $1.221$ $1.434$ $1.052$ $1.156$ $1.124$ $1.268$ $1.973$ $2.061$ $1.853$ $1.899$ $2.043$ $2.058$
+PROD_OF_FRO $45.667$ $52.105$ $159.109$ $210.065$ $4.538$ $5.582$ $42.519$ $50.772$ $129.111$ $141.794$ $6.510$ $7.677$ $48.816$ $53.437$
+PROD_OF_FRO_OVER_MARGIN $45.569$ $52.134$ $158.752$ $210.292$ $4.520$ $5.538$ $42.487$ $50.732$ $129.093$ $141.846$ $6.391$ $7.530$ $48.650$ $53.537$
+SPEC_ORIG_MAIN $66.789$ $72.385$ $177.665$ $232.486$ $5.294$ $6.130$ $55.105$ $64.848$ $195.785$ $209.989$ $7.909$ $8.296$ $78.473$ $79.923$
+PROD_OF_SPEC $67.284$ $72.701$ $177.627$ $232.516$ $5.420$ $6.298$ $55.458$ $64.973$ $196.634$ $210.576$ $8.272$ $8.678$ $79.109$ $80.429$
+PROD_OF_SPEC_OVER_MARGIN $67.196$ $72.740$ $177.751$ $232.603$ $5.401$ $6.235$ $55.227$ $64.942$ $196.343$ $210.513$ $8.133$ $8.548$ $79.165$ $80.539$
+SPEC_INIT_MAIN $67.235$ $72.897$ $177.757$ $232.667$ $5.432$ $6.237$ $55.198$ $64.927$ $197.145$ $210.896$ $8.288$ $8.751$ $79.271$ $80.867$
+Top numbers: CMS; bottom numbers: eCMS.
+Table 3: CMS and eCMS summary $\delta=0.05$
+Measure $\mathrm{CMS}^{\mathrm{med}}_{\delta}/\mathrm{eCMS}^{\mathrm{med}}_{\delta}$ C10/DN121 C10/NiN C10/RN50 FMN/DN121 FMN/NiN FMN/RN50
+PARAMS $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$
+PACBAYES_MAG_ORIG $0.096$ $0.096$ $0.117$ $0.131$ $0.056$ $0.054$ $0.075$ $0.076$ $0.110$ $0.113$ $0.082$ $0.079$ $0.120$ $0.119$
+PACBAYES_MAG_INIT $0.136$ $0.144$ $0.133$ $0.143$ $0.067$ $0.064$ $0.139$ $0.145$ $0.286$ $0.289$ $0.073$ $0.068$ $0.302$ $0.290$
+FRO_OVER_SPEC $0.188$ $0.198$ $0.138$ $0.162$ $0.143$ $0.143$ $0.095$ $0.111$ $0.369$ $0.389$ $0.233$ $0.234$ $0.561$ $0.569$
+INVERSE_MARGIN $0.246$ $0.210$ $0.228$ $0.191$ $0.167$ $0.155$ $0.265$ $0.229$ $0.323$ $0.297$ $0.145$ $0.127$ $0.364$ $0.326$
+PACBAYES_MAG_FLATNESS $0.299$ $0.311$ $0.395$ $0.442$ $0.221$ $0.204$ $0.310$ $0.347$ $0.288$ $0.275$ $0.170$ $0.141$ $0.411$ $0.381$
+PACBAYES_FLATNESS $0.627$ $0.628$ $0.448$ $0.478$ $0.539$ $0.540$ $0.392$ $0.426$ $0.715$ $0.717$ $0.754$ $0.738$ $0.796$ $0.782$
+PACBAYES_ORIG $0.558$ $0.665$ $0.550$ $0.655$ $0.972$ $0.989$ $0.566$ $0.676$ $0.473$ $0.461$ $0.751$ $0.744$ $0.505$ $0.470$
+SUM_OF_FRO_OVER_MARGIN $0.781$ $0.862$ $0.622$ $0.757$ $1.033$ $1.132$ $0.763$ $0.917$ $0.498$ $0.545$ $0.911$ $1.070$ $0.799$ $0.807$
+SUM_OF_FRO $0.785$ $0.867$ $0.623$ $0.757$ $1.045$ $1.141$ $0.765$ $0.918$ $0.501$ $0.547$ $0.929$ $1.090$ $0.805$ $0.816$
+PARAM_NORM $0.955$ $1.097$ $0.985$ $1.164$ $1.359$ $1.445$ $0.854$ $1.030$ $0.702$ $0.774$ $1.281$ $1.321$ $0.925$ $0.975$
+PATH_NORM_OVER_MARGIN $1.288$ $1.218$ $1.220$ $1.134$ $0.445$ $0.403$ $1.356$ $1.303$ $3.325$ $3.351$ $0.593$ $0.555$ $4.025$ $4.024$
+SUM_OF_SPEC_OVER_MARGIN $1.115$ $1.235$ $0.744$ $0.856$ $1.216$ $1.252$ $1.069$ $1.251$ $0.819$ $0.845$ $1.161$ $1.219$ $1.468$ $1.488$
+SUM_OF_SPEC $1.124$ $1.245$ $0.744$ $0.857$ $1.227$ $1.269$ $1.068$ $1.253$ $0.819$ $0.848$ $1.180$ $1.238$ $1.473$ $1.489$
+PACBAYES_INIT $1.241$ $1.254$ $0.973$ $1.012$ $1.227$ $1.257$ $0.870$ $0.917$ $1.777$ $1.803$ $1.256$ $1.251$ $1.864$ $1.849$
+PATH_NORM $1.367$ $1.296$ $1.322$ $1.242$ $0.556$ $0.520$ $1.413$ $1.349$ $3.133$ $3.149$ $0.632$ $0.590$ $3.930$ $3.921$
+DIST_SPEC_INIT $1.659$ $1.708$ $1.442$ $1.561$ $1.578$ $1.605$ $1.417$ $1.549$ $2.577$ $2.656$ $1.740$ $1.810$ $2.778$ $2.777$
+FRO_DIST $1.763$ $1.807$ $1.387$ $1.453$ $1.670$ $1.713$ $1.269$ $1.386$ $2.500$ $2.535$ $1.856$ $1.901$ $2.646$ $2.653$
+PROD_OF_FRO_OVER_MARGIN $45.293$ $50.008$ $150.478$ $183.171$ $7.228$ $7.921$ $44.252$ $53.202$ $120.482$ $131.962$ $6.377$ $7.487$ $46.334$ $46.814$
+PROD_OF_FRO $45.538$ $50.283$ $150.676$ $183.231$ $7.314$ $7.989$ $44.385$ $53.258$ $121.211$ $132.342$ $6.503$ $7.627$ $46.690$ $47.308$
+SPEC_ORIG_MAIN $73.068$ $78.684$ $179.892$ $206.998$ $8.362$ $8.650$ $61.769$ $71.864$ $197.573$ $204.030$ $7.901$ $8.281$ $84.367$ $85.503$
+PROD_OF_SPEC_OVER_MARGIN $73.576$ $79.426$ $180.133$ $207.233$ $8.512$ $8.765$ $62.025$ $72.539$ $198.114$ $204.537$ $8.128$ $8.533$ $85.126$ $86.313$
+PROD_OF_SPEC $73.690$ $79.493$ $180.019$ $207.366$ $8.589$ $8.882$ $61.970$ $72.646$ $198.217$ $205.168$ $8.262$ $8.664$ $85.409$ $86.340$
+SPEC_INIT_MAIN $73.884$ $79.766$ $180.209$ $207.184$ $8.527$ $8.844$ $61.985$ $72.541$ $199.461$ $206.594$ $8.284$ $8.740$ $85.783$ $86.990$
+Top numbers: CMS; bottom numbers: eCMS.
+A note on scope.
+Our quantitative fragility scores target a single desideratum: when test error does not change, a credible generalization measure should remain stable.
+Scoring well on CMS/eCMS does not imply that a measure is “correct” or sufficient.
+For example, VC dimension/parameter count is widely regarded as a poor complexity measure for modern deep networks, yet it is not fragile in our sense (cf.
+PARAMS rows).
+Likewise, the ML‑PACBayes bound discussed in the main text (§F) is not fragile largely because it is agnostic to hyperparameter changes.
+Stability is necessary for utility, but it is not, by itself, a full explanation of generalization.
+Appendix F Post-mortem vs. ML-PACBayes
 We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of Valle-Pérez & Louis (2020).
@@
 All figure numbers below refer to Valle-Pérez & Louis (2020).
-Definition E.1 (Marginal-likelihood PAC–Bayes bound).
+Definition F.1 (Marginal-likelihood PAC–Bayes bound).
 Consider binary classification with data distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$ and a hypothesis space $\mathcal{H}$ of functions $h:\mathcal{X}\to\{0,1\}$.
@@
 $\ln\!\bigl(1-\varepsilon(h)\bigr)\;<\;\frac{\ln\!\tfrac{1}{P(C(S))}+\ln n+\ln\!\tfrac{1}{\delta}+\ln\!\tfrac{1}{\gamma}}{\,n-1\,}.$ (24)
-Remark E.2 (binary vs. multiclass).
+Remark F.2 (binary vs. multiclass).
 We state the bound for binary classification for clarity; the function‑space marginal‑likelihood perspective extends to multiclass via standard reductions (e.g., one‑vs‑rest) or multiclass likelihoods.
@@
 We view this invariance as a virtue for a predictor; the bound is not designed to explain differences caused by optimizer choice and should not be judged on that criterion.
-These properties are consistent with our optimizer‑swap and pixel‑permutation stress tests (Apps. B and F) and with our scale‑symmetry Exp++ dial, which only alters parameter scale (App. G).
+These properties are consistent with our optimizer‑swap and pixel‑permutation stress tests (Apps. B and G) and with our scale‑symmetry Exp++ dial, which only alters parameter scale (App. H).
 Why does the function‑space route outperform post‑mortems?
@@
 Closing this gap—by designing reparameterization‑invariant, data‑aware post‑mortem diagnostics that inherit the Occam flavour of marginal likelihood—remains open.
-Appendix F Stress-testing generalization measures with pixel permutations
+Appendix G Stress-testing generalization measures with pixel permutations
 A useful generalization measure should pass two basic stress tests: (i) it should be insensitive to symmetry-preserving transformations that do not change the intrinsic task, yet (ii) sensitive when task-relevant information is destroyed.
@@
 best accuracy “Acc”) to expose hyperparameter sensitivity.
-Results are summarized in Tables 1 and 2.
+Results are summarized in Tables 4 and 5.
 FCN (permutation symmetry holds).
@@
   • CE: best checkpoint by cross-entropy; Acc: best checkpoint by accuracy.
-Appendix G Exp++ in scale-invariant nets: protocol and results
+Appendix H Exp++ in scale-invariant nets: protocol and results
 We instantiate the symmetry in a fully connected, scale‑invariant network (normalization after each hidden linear layer; the final linear readout is frozen), trained on MNIST with SGD + momentum.

### v3

@@
 Abstract
-A wide variety of generalization measures have been applied to deep neural networks (DNNs).
-Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends.
-In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behavior.
-For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalization measures like the path norm.
+In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks—are fragile: small training modifications that barely affect the performance of the underlying deep neural network can substantially change a measure’s value, trend, or scaling behavior.
+For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants, can reverse the slope of a learning curve in widely used generalization measures such as the path norm.
 We also identify subtler forms of fragility.
@@
 This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behavior, in learning curves, but which is not a post-mortem measure.
-Beyond demonstrating that many bounds—such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC-Bayes surrogates—are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
+Beyond demonstrating that many post-mortem bounds are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
 Machine Learning, ICML
-1 Introduction
+1 Introduction to the position statement
 Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning.
@@
 For deep networks the same aspiration has driven a profusion of bounds.
-While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs).
+While tight bounds have proven elusive, with some exceptions (dziugaite2017computing; prezortiz2020tighter; valleprez2020generalization; lotfi2022pacb), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs).
 These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds.
 Large-scale comparisons have compared the behavior of these measures, albeit often only at the sign level, e.g.
-does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
-A notable property of DNNs is that to first order, their generalization performance is robust to modest changes in network size, hyperparameters, optimisers, or stopping criteria.
-This raises a natural question: are generalization measures themselves equally robust to such changes?
+does changing a setting produce the same change in the underlying DNN and a generalization measure with the same sign (jiang2019fantastic; dziugaite2020in).
+A notable property of DNNs is that to first order, generalization performance is robust to modest changes in network size, hyperparameters, optimizers, or stopping criteria.
+This raises a natural question: do generalization measures share this robustness?
 If a generalization measure captures the basic underlying factors that drive DNN performance, one would expect it to exhibit a similar degree of robustness.
-The main finding of this paper is that many of the post-mortem bounds studied in the literature are fragile: small, ostensibly irrelevant training tweaks that barely affect the underlying DNN, can flip a measure’s shape, trend, or scale.
-To test these fragilities systematically, we evaluate measures under controlled perturbations that hold data and architecture, fixed while nudging a single training knob.
+The main position this paper argues for is that: Many of the post-mortem bounds studied in the literature are fragile: small training tweaks that barely affect the underlying DNN performance can drastically change a measure’s behaviour.
+Investigators should carefully audit their generalization measures for fragility, using the kinds of tests we propose in this paper.
+To test these fragilities systematically, we develop a fragility audit that evaluates measures under controlled perturbations that hold data and architecture fixed while nudging a single training knob.
 Three stressors structure the audit in the main text: learning‑curve behavior as sample size increases (§3); temporal behavior, including when the model has interpolated (§4); and responses to data complexity via label noise and dataset swaps (§5).
-In the supplementary material, we show more tests of fragility for a much wider set of measures.
-In contrast to the post-mortem bounds, we show that a function-based marginal-likelihood PAC-Bayes bound successfully tracks dataset complexity and learning-curve scaling (Valle-Pérez & Louis, 2020).
+§6 reveals that the tightest existing bound fails across both its bias and complexity components.
+In the supplementary material, we test a much wider set of measures and develop an expanded fragility audit.
+In contrast to the post-mortem bounds, we show that a function-based marginal-likelihood PAC-Bayes bound successfully tracks dataset complexity and learning-curve scaling (valleprez2020generalization).
 This bound is based on a simple measure of inductive bias, the probability that a random set of weights produces a function that fits the training data set.
-Because it operates in function space and does not depend explicitly on training dynamics, this bound is inherently insensitive to the choice of optimiser or most training hyperparameters.
-This insensitivity is both a strength, making the bound robust rather than fragile, and a limitation, as it prevents the bound from reflecting how training procedures influence generalization, including effects such as feature learning.
+Because it operates in function space and does not depend explicitly on training dynamics, this bound is inherently insensitive to the choice of optimizer or most training hyperparameters.
+This insensitivity is both a strength, making the bound robust rather than fragile, and a limitation, as it prevents the bound from reflecting how training procedures influence generalization, including mechanisms such as feature learning.
 Nonetheless, the strong performance of this relatively simple approach suggests that developing more sophisticated function-space generalization measures could be a fruitful direction for future research.
 Finally, we sketch a few potential causes of fragility, in the hope they can lead to new insights.
-We show that some of the qualitative failures of norm-based bounds with sample size appear to track a recent prediction of qualitative changes in the scaling of norms in the simpler setting of linear regression (Zhang & Louis, 2025).
+We show that some of the qualitative failures of norm-based bounds with sample size appear to track a recent prediction of qualitative changes in how norms scale with data set size in the simpler setting of linear regression (zhang2025closed).
 We also exploit a symmetry in scale‑invariant networks and prove a non‑asymptotic equivalence between fixed learning‑rate with fixed weight decay and a matched run with exponentially increasing learning rate and time‑varying weight decay; the two training procedures compute the same predictor at every iterate (§7; cf.
-Li & Arora, 2019).
+li2019an).
 Under this invariance lever, magnitude‑sensitive post‑mortem measures can inflate by orders of magnitude while test error remains flat.
 This suggests that generalization measures should carefully track key invariances present in the underlying DNNs.
-The widespread fragility we observe suggests that some generalization measures may be showing correlation in some limited settings that may not be related to causation.
-Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
-Our contribution:
-  • We formulate a compact fragility audit for generalization measures that targets training-hyperparameter stability, post‑interpolation temporal behavior, and data‑complexity response.
-  • We provide systematic evidence that popular post‑mortem measures—path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates—change their qualitative predictions under mild optimiser or hyperparameter tweaks even when the underlying DNN accuracy is stable.
-  • We prove an equivalence for scale‑invariant networks that matches fixed schedules to exponentially increasing learning‑rate schedules with time‑varying weight decay, yielding a controlled invariance lever that isolates function from parameter scale and exposes magnitude sensitivity.
+We issue a call to action: Further work to understand the causes of fragility is essential. Otherwise it is hard to assess whether a generalization measures reflects a true causal relationship or not.
+Our contribution: We (1) formulate a stress-test protocol for hyperparameter stability, temporal drift, and data complexity; (2) demonstrate that popular post-mortem measures are fragile, often reversing trends or failing to track data difficulty under benign tweaks that leave the underlying predictor unchanged.; (3) introduce a quantitative score for measure instability; and (4) prove a non-asymptotic equivalence between fixed and exponential LR schedules for scale-invariant nets, exposing flaws in magnitude-sensitive measures.
 2 Related work
@@
 From classical capacity to modern practice.
-Early theory framed capacity via VC and Rademacher analyses (Mendelson, 2002; Anthony & Bartlett, 2002; Bartlett & Mendelson, 2001; Koltchinskii, 2001).
-As overparameterization became the norm, evidence accumulated that uniform-convergence explanations often miss the mark: networks can fit random labels yet generalize on natural data (Zhang et al., 2016), some bounds fail to track risk (Nagarajan & Kolter, 2019), and benign overfitting can arise even at interpolation (Bartlett et al., 2020).
+Early theory framed capacity via VC and Rademacher analyses (vapnik1998statistical; anthony1999neural; bartlett2001rademacher; koltchinskii2001rademacher).
+As overparameterization became the norm, evidence accumulated that uniform-convergence explanations often miss the mark: networks can fit random labels yet generalize on natural data (zhang2016understandingc), some bounds fail to track risk (nagarajan2019uniform), and benign overfitting can arise even at interpolation (bartlett2020benign).
 The field responded with diagnostics that foreground invariances, algorithmic dependence, and data interactions.
-2.1 Capacity-oriented diagnostics: norms, margins, distance from initialization
-The first works in this subfield translated classical notions into deep settings.
-Spectrally-normalized margin bounds tied test error to Lipschitz-like control and margins, making scale explicit (Bartlett et al., 2017).
-Path- and norm-based views clarified how depth and weight scales shape effective complexity (Neyshabur et al., 2015, 2019).
-Empirically, modeling the distribution of margins—not just the minimum—improves predictiveness across trained families (Jiang et al., 2018; Novak et al., 2018), and sample complexity can depend on norms rather than width (Golowich et al., 2018).
-In practice, enforcing Lipschitz continuity often improves out-of-sample performance (Gouk et al., 2020; Yoshida & Miyato, 2017).
-Distance-from-initialization and movement-from-pretraining offer reference-aware, width-robust complexity surrogates with both empirical support and bounds (Li et al., 2018; Zhou et al., 2021; Neyshabur et al., 2017a), while optimization dynamics link large margins to implicit bias in separable regimes (Soudry et al., 2018).
-2.2 Geometry-oriented diagnostics: flatness and sharpness
+Capacity-oriented diagnostics: norms, margins, distance from initialization.
+Early work translated classical notions into deep settings.
+Spectrally-normalized margin bounds tied test error to Lipschitz-like control and margins, making scale explicit (bartlett2017spectrallyb).
+Path- and norm-based views clarified how depth and weight scales shape effective complexity (neyshabur2015norm; neyshabur2019theb).
+Empirically, modeling the distribution of margins, not just the minimum, improves predictiveness across trained families (jiang2018predicting; novak2018sensitivity), and sample complexity can depend on norms rather than width (golowich2018size).
+In practice, enforcing Lipschitz continuity often improves out-of-sample performance (gouk2020regularisation; yoshida2017spectral).
+Distance-from-initialization and movement-from-pretraining offer reference-aware, width-robust complexity surrogates with both empirical support and bounds (li2018on; zhou2021understanding; neyshabur2017exploring), while optimization dynamics link large margins to implicit bias in separable regimes (soudry2018the).
+Geometry-oriented diagnostics: flatness and sharpness.
 A parallel line approached generalization through loss-landscape geometry.
-The “flat minima generalize better” intuition predates deep learning (Hochreiter & Schmidhuber, 1997) and resurfaced when large-batch training was observed to converge to sharper minima with worse test error (Keskar et al., 2017).
-Because raw sharpness is parameterization-dependent, normalized and magnitude-aware definitions were proposed and shown to correlate more robustly with test error (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022).
-These ideas also shaped algorithms: SAM explicitly optimizes a local worst-case neighborhood and typically improves accuracy; adaptive variants and ablations probe when and why it helps (Foret et al., 2020; Kwon et al., 2021; Andriushchenko & Flammarion, 2022).
-Independent probes—weight averaging, landscape visualizations, and training-dynamics analyses—add evidence that broader valleys often accompany better generalization (Izmailov et al., 2018; Li et al., 2017; Cohen et al., 2021).
-Stochastic optimization theory offers a mechanism: noise scale and heavy-tailed gradient noise can bias learning toward flatter regions (Smith et al., 2017; Jastrzebski et al., 2017; Simsekli et al., 2019; McCandlish et al., 2018).
-2.3 Algorithm-aware certificates: PAC-Bayes as operational measures
-Partly in response to limits of uniform convergence, PAC-Bayesian analysis made algorithm dependence explicit by trading empirical risk against a posterior–prior divergence (McAllester, 1999; Seeger, 2002; Catoni, 2007; Langford & Shawe-Taylor, 2002).
-The framework became operational when non-vacuous deep-network certificates were obtained by optimizing stochastic posteriors (Dziugaite & Roy, 2017), or simply working with the GP prior (Valle-Pérez et al., 2018; Zhang et al., 2021).
-Subsequent work tightened certificates with data-aware priors (Dziugaite, 2018; Pérez‐Ortiz et al., 2020), and with compression-flavored posteriors that reduce effective description length (Arora et al., 2018a; Lotfi et al., 2022).
-Treating PAC-Bayes itself as a training objective yields models that are both accurate and tightly certified, while extensions broaden the scope to adversarial risk and fast/mixed-rate regimes (Rodríguez-Gálvez et al., 2024).
-However, one worry is that while these optimised measures may provide quantitatively tighter bounds, they do so at the expense of qualitative insight (Arora et al., 2018a).
+The “flat minima generalize better” intuition predates deep learning (hochreiter1997flatb) and resurfaced when large-batch training was observed to converge to sharper minima with worse test error (keskar2016on).
+Because raw sharpness is parameterization-dependent, normalized and magnitude-aware definitions were proposed and shown to correlate more robustly with test error (dinh2017sharpb; tsuzuku2020normalized; kim2022scale; jang2022reparametrization).
+These ideas also shaped algorithms: SAM explicitly optimizes a local worst-case neighborhood and typically improves accuracy; adaptive variants probe when and why it helps (foret2020sharpnessb; kwon2021asam; zela2022towards).
+Independent probes—weight averaging, landscape visualizations, and training-dynamics analyses—add evidence that broader valleys often accompany better generalization (izmailov2018averagingb; li2017visualizing; cohen2021gradient).
+Stochastic optimization theory offers a mechanism: noise scale and heavy-tailed gradient noise can bias learning toward flatter regions (smith2017donb; jastrzbski2017three; simsekli2019tail; mccandlish2018an).
+Algorithm-aware certificates: PAC-Bayes as operational measures.
+Partly in response to limits of uniform convergence, PAC-Bayesian analysis made algorithm dependence explicit by trading empirical risk against a posterior–prior divergence (mcallester1999someb; seeger2002pac; catoni2007pac; langford2002pac).
+The framework became operational when non-vacuous deep-network certificates were obtained by optimizing stochastic posteriors (dziugaite2017computing), or simply working with the GP prior (valleprez2018deepb; zhang2021why).
+Subsequent work tightened certificates with data-aware priors (dziugaite2018data; prezortiz2020tighter), and with compression-flavored posteriors that reduce effective description length (arora2018strongerc; lotfi2022pacb).
+Treating PAC-Bayes itself as a training objective yields models that are both accurate and tightly certified, while extensions broaden the scope to adversarial risk and fast/mixed-rate regimes (rodrguezglvez2024mixed).
+However, one worry is that while these optimised measures may provide quantitatively tighter bounds, they do so at the expense of qualitative insight (arora2018strongerc).
 Other perspectives.
 Complementary lenses provide boundary conditions any credible measure should respect.
-Algorithmic stability ties generalization to insensitivity under data perturbations (Bousquet & Elisseeff, 2002; Hardt et al., 2015); information-theoretic analyses bound excess risk by the information a learning rule extracts from the sample (Russo & Zou, 2016; Xu & Raginsky, 2017).
-Description-length and compression viewpoints link compressibility to generalization (Blier & Ollivier, 2018; Taiji Suzuki, 2020).
-Linearized regimes (NTK/GP) delineate when very wide nets behave like their kernel limits (Jacot et al., 2018; Lee et al., 2020b).
-Finally, double-descent phenomena and scaling laws offer external checks on diagnostics (Belkin et al., 2019; Kaplan et al., 2020).
+Algorithmic stability ties generalization to insensitivity under data perturbations (bousquet2002stability; hardt2016train); information-theoretic analyses bound excess risk by the information a learning rule extracts from the sample (russo2016controlling; xu2017information).
+Description-length and compression viewpoints link compressibility to generalization (blier2018the; suzuki2020compression).
+Linearized regimes (NTK/GP) delineate when very wide nets behave like their kernel limits (jacot2018neuralb; lee2019wide).
+Finally, double-descent phenomena and scaling laws offer external checks on diagnostics (belkin2019reconciling; kaplan2020scaling).
 Meta-evaluations and synthesis.
-Large-scale comparisons underscore that no single candidate explains generalization under all interventions (Jiang et al., 2019b; Dziugaite et al., 2020b), motivating a measures-as-diagnostics mindset.
-A recent critique shifts the lens from predictiveness to tightness, showing that uniformly tight bounds are out of reach in overparameterized settings and clarifying what such tightness could mean (Gastpar et al., 2023).
+Large-scale comparisons underscore that no single candidate explains generalization under all interventions (jiang2019fantastic; dziugaite2020in), motivating a measures-as-diagnostics mindset.
+A recent critique shifts the lens from predictiveness to tightness, showing that uniformly tight bounds are out of reach in overparameterized settings and clarifying what such tightness could mean (gastpar2023fantastic).
 Our focus is complementary: rather than tightness per se, we study fragility—how otherwise informative measures can fail under innocuous hyperparameter changes that leave underlying DNN performance essentially unchanged.
@@
 The path norm (definition in App.
-A) is often regarded as the strongest norm-based proxy for modern ReLU networks: it respects layerwise rescalings and has supported some of the sharpest norm-style capacity statements and practical diagnostics (Neyshabur et al., 2015, 2017b; Gonon et al., 2023).
+A) is often regarded as the strongest norm-based proxy for modern ReLU networks: it respects layerwise rescalings and has supported some of the sharpest norm-style capacity statements and practical diagnostics (neyshabur2015norm; neyshabur2017a; gonon2023a).
 Precisely because it is a strong candidate, it is a natural place to look for fragility.
@@
 1 we show a triptych for ResNet‑50 on FashionMNIST, nudging exactly one hyperparameter at a time, either the optimizer or the learning rate (no early stopping), while keeping the data, model, and pipeline fixed.
-Each curve shows test error (black), the raw path norm (blue), and the path norm divided by the empirical margin (orange); axes are logarithmic and error bars reflect seed variability.
+Each curve shows test error (black), the raw path norm (blue), and the path norm divided by the empirical margin (orange).
+In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower.
+Both drop faster than the underlying ResNet‑50 test error.
+Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior.
+A minor adjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows.
+These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
 Figure 1: Path‑norm learning curves under small training‑pipeline changes (no early stopping).
@@
 Training set sizes $n$ range from $10^{2}$ to the full FashionMNIST training set ($5\times 10^{4}$ samples).
-In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower.
-Both drop faster than the underlying ResNet‑50 test error.
-Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior.
-A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows.
-These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
 The contrasting behavior of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression.
 In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer.
-More precisely, with $p\in[1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if $r\leq 2(p-1)$  (Zhang & Louis, 2025).
+More precisely, with $p\in(1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if r>2(p1)r>2(p-1) (zhang2025closed).
 This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm.
-This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (Zhang & Louis, 2025).
-For the more complex setup of path-norm for ResNet‑50 on FashinMNIST in Fig.
+This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (zhang2025closed).
+For the more complex setup of path-norm for ResNet‑50 on FashionMNIST in Fig.
 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
-See App. A for precise definitions for the measures we consider as well as further experiments, including the same stress test across Frobenius, margin, spectral, PAC‑Bayes, and VC‑style proxies.
+See App.
+A for precise definitions for the measures we consider as well as further experiments, including the same stress test across Frobenius, margin, spectral, PAC‑Bayes, and VC‑style proxies [^1] All experiments in this work are fully reproducible with minimum hardware requirements.
+The anonymized code can be found at https://anonymous.4open.science/r/Prior_bound-8A6C/..
 4 Temporal behavior fragility
@@
 Adam tends to amplify this scale drift, while SGD+momentum tempers it, yielding opposite signs for the post‑interpolation slope.
-The lesson is that weight–norm–based measures—and margin‑normalized variants that co‑move with them—are not monotone indicators of optimization progress: an upward trend can be optimizer‑driven scale inflation; a flat or gently declining trend need not signal stagnation.
-For practice, we recommend reporting $T_{\mathrm{int}}$ and the post‑interpolation slope $s\coloneqq\tfrac{d\log\mathrm{Measure}}{d\log t}\big|_{t>T_{\mathrm{int}}}$ alongside any temporal plot; overlay optimizers on the same axes; and compare snapshots at matched milestones (e.g., the first $100\%$‑accuracy crossing) or with early‑stopped checkpoints.
-A useful stress test is hysteresis: resume a checkpoint taken just after $T_{\mathrm{int}}$ with a different optimizer and check whether the measure’s drift changes sign while test error holds steady.
-These checks make temporal fragility visible and keep optimizer‑induced motion from being mistaken for learning.
+The lesson is that weight-norm-based measures—and margin-normalized variants that co-move with them—are not monotone indicators of optimization progress: an upward trend can be optimizer‑driven scale inflation; a flat or gently declining trend need not signal stagnation.
 We refer the readers to Appendix B for experimental results on the temporal behaviors of more generalization measures.
@@
 There has been substantial recent interest in the study of scaling laws for deep neural networks (DNNs).
-These laws describe how generalization performance varies as a function of model size, training sample size, or computational cost (see, e.g., (Hestness et al., 2017; Kaplan et al., 2020; Spigler et al., 2019; Hoffmann et al., 2022; Valle-Pérez & Louis, 2020; Bahri et al., 2021; Nam et al., 2024; Goring et al., 2025) and references therein).
+These laws describe how generalization performance varies as a function of model size, training sample size, or computational cost (see, e.g., (hestness2017deep; kaplan2020scaling; spigler2019asymptotic; hoffmann2022training; valleprez2020generalization; bahri2024explaining; nam2024visualising; goring2025feature) and references therein).
 Here, we observe that the ML-PACBayes bound does quite well at capturing the scaling behavior with the training sample size.
-However, it should be kept in mind that because this bound is evaluated with a Gaussian process (GP) method, it cannot capture feature-learning, which can play a crucial role in learning-curve scaling for certain datasets (Nam et al., 2024; Goring et al., 2025).
+However, it should be kept in mind that because this bound is evaluated with a Gaussian process (GP) method, it cannot capture feature-learning, which can play a crucial role in learning-curve scaling for certain datasets (nam2024visualising; goring2025feature).
 Figure 4: Fixed architecture, varying dataset..
@@
 Changing from ADAM to SGD with momentum leads to a small increase in the overall scale of the PACBAYES_ORIG learning curves (solid lines).
-However, for both optimisers, these learning curves don’t distinguish MNIST, FashionMNIST, and CIFAR10, missing both the data‑complexity ordering and the changing slopes visible in the underlying learning curves (dashed lines).
+However, for both optimizers, these learning curves don’t distinguish MNIST, FashionMNIST, and CIFAR10, missing both the data‑complexity ordering and the changing slopes visible in the underlying learning curves (dashed lines).
 Right: ML‑PACBayes (marginal‑likelihood PAC–Bayes) is tight and preserves the correct dataset ordering and reproduces the trends in the slope; for technical reasons, this panel is computed on the binarized versions of the three datasets.
@@
 Again, we find that the ML-PACBayes function‑space predictor tracks the expected invariances while several post‑mortem diagnostics do not.
+6 Fragility of tightness: the case of compression bounds
+Recent work by lotfi2022pacb proposes PAC-Bayes compression bounds that achieve state-of-the-art tightness.
+They argue that this tightness allows them to “explain” generalization.
+While this approach is theoretically stimulating, our audit suggests that numerical tightness does not necessarily imply a correct causal explanation of the generalization phenomenon.
+We identify two specific fragilities.
+(a) CIFAR-10 (ID 3500) Bounds
+(b) MNIST (ID 1000) Bounds
+(c) KL Divergence (CIFAR vs MNIST)
+Figure 5: Decomposing PAC-Bayes Compression Bounds.
+Left & Middle: Learning curves for bounds and error rates.
+The “quantized training error” (dashed brown) is large and tracks the test error.
+Right: The Raw KL divergence (complexity term) remains essentially flat for both datasets across sample sizes ($n$).
+Additional results for higher intrinsic dimensions are provided in Appendix I.
+Bias dominance over complexity.
+In standard learning theory, bounds typically trade off a small training error against a complexity penalty.
+However, as shown in the left and middle panels of Figure 5, the “quantized training error” (dashed brown line)—which serves as the empirical risk term in the bound—is notably large.
+Crucially, as the sample size $n$ increases, this quantized training error does not vanish but instead grows to track the quantized test error.
+This suggests that the final bound’s tightness is driven primarily by this large bias term (the quantization error) rather than by the KL-divergence complexity term measuring the neural network’s capacity.
+In practice, modern DNNs often achieve near-zero training error; a bound that relies on a large proxy training error to achieve tightness is arguably measuring the quality of the compression scheme rather than the generalization mechanism of the original network.
+Insensitivity of the complexity term.
+A generalization measure should ideally reflect how the difficulty of the learning problem changes with data availability.
+However, the right panel of Figure 5 demonstrates that the KL divergence—the core measure of complexity in the PAC-Bayes formulation—remains almost perfectly flat with respect to training set size $n$.
+This flatness persists across different datasets (CIFAR-10 vs.
+MNIST) and intrinsic dimensions (see Appendix I for results with $\text{ID}=10000$).
+This implies that the complexity term fails to capture the scaling of learning difficulty, rendering the bound insensitive to the actual information gain from increasing data.
 7 Scale-invariant network and exponential learning rate schedule
-Normalization layers make many modern networks effectively scale‑invariant, and this symmetry underlies the “exponentially increasing learning‑rate” idea of Li & Arora (2019): in such models, training with fixed learning rate (LR), weight decay (WD), and momentum can be matched—in function space—by a run with an exponentially increasing LR and no WD (plus a small momentum correction across phases).
-Their analysis shows how LR and WD can be “folded into” an exponential schedule and documents successful experiments under this equivalence.
+Normalization layers render many modern networks effectively scale-invariant, motivating the exponentially increasing learning-rate scheme of li2019an.
+In Appendix C, we prove a non-asymptotic equivalence: training with fixed learning rate (LR) and weight decay (WD) is exactly matched in function space by an exponentially increasing LR with time-varying WD.
 We refer to this multiplicative control parameter $\alpha$ as the Exp++ factor.
-We use the same symmetry but a different equivalence.
-Instead of removing WD, our theorem pairs an exponentially increasing LR with a time‑varying WD so that every iterate computes the same predictor as a fixed‑LR/fixed‑WD baseline.
-The mapping is explicit and non‑asymptotic: it gives closed‑form schedules for $\tilde{\eta}_{t}$ and $\tilde{\lambda}_{t}$, and an admissible range for the multiplicative factor $\alpha$ via the interval $\mathcal{I}$.
-Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures.
-Empirical details and results are deferred to Appendix H.
-Definition 6.1 (Scale invariant neural networks).
-Consider a parameterized neural network $f(\boldsymbol{\theta})$. We say $f$ is scale invariant if
-$\forall c\in\mathbb{R}^{+},f(\boldsymbol{\theta})=f(c\boldsymbol{\theta})$ (1)
-Theorem 6.2 (Equivalence of schedules in scale-invariant nets).
-[…]
-$\Delta_{\lambda}$ $\coloneqq(1-\gamma)^{2}-2(1+\gamma)\lambda\eta_{0}+(\lambda\eta_{0})^{2},$ (2)
-$\Xi(\alpha)$ $\coloneqq\frac{\alpha^{2}-\alpha(1-\lambda\eta_{0}+\gamma)+\gamma}{\eta_{0}}.$ (3)
-Define the interval endpoints
-$\alpha_{L}$ $\coloneqq\frac{\gamma}{\,1-\lambda\eta_{0}+\gamma\,}$, (4)
-$\alpha_{-}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}-\sqrt{\Delta_{\lambda}}}{2},$ (5)
-$\alpha_{+}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}+\sqrt{\Delta_{\lambda}}}{2},$ (6)
-and set
-$\mathcal{I}\coloneqq(\alpha_{L},\ \alpha_{-}]\ \cup\ [\alpha_{+},\,1).$ (7)
-Consider the two updates (with $\theta_{-1}=\theta_{0}$, $\tilde{\theta}_{0}=\theta_{0}$, $\tilde{\theta}_{-1}=\alpha\,\theta_{-1}$):
-$\frac{\theta_{t}-\theta_{t-1}}{\eta_{0}}$ $=\frac{\gamma(\theta_{t-1}-\theta_{t-2})}{\eta_{0}}-\nabla_{\theta}\!\Big(L(\theta_{t-1})+\tfrac{\lambda}{2}\|\theta_{t-1}\|_{2}^{2}\Big),$ (A)
-$\frac{\tilde{\theta}_{t}-\tilde{\theta}_{t-1}}{\tilde{\eta}_{t}}$ $=\frac{\gamma(\tilde{\theta}_{t-1}-\tilde{\theta}_{t-2})}{\tilde{\eta}_{t-1}}-\nabla_{\theta}\!\Big(L(\tilde{\theta}_{t-1})+\tfrac{\tilde{\lambda}_{t}}{2}\|\tilde{\theta}_{t-1}\|_{2}^{2}\Big).$ (B)
-If $\alpha\in\mathcal{I}$, then the schedules
-$\tilde{\eta}_{t}$ $=\eta_{0}\,\alpha^{-2t-1},$ (8)
-$\tilde{\lambda}_{t}$ $=\Xi(\alpha)\,\alpha^{2t-1}+\frac{\gamma(1-\alpha)}{\eta_{0}\,\alpha}\,\mathbbm{1}\{t=0\}$ (9)
-ensure $\tilde{\theta}_{t}=\alpha^{-t}\theta_{t}$ for all $t\geq 0$; hence (A) and (B) generate identical functions.
-Remark 6.3 (Parameter range and nonnegativity).
-Assume
-$\lambda\eta_{0}$ $\leq(1-\sqrt{\gamma})^{2},$ (10)
-$\alpha_{L}$ $<\alpha_{-}\quad\text{(equivalently, $\ \Delta_{\lambda}\geq 0$)}.$ (11)
-Then the interval $\mathcal{I}$ above is nonempty and the square‑root term is real; both conditions hold for common hyperparameters in practice.
-Consequences.
-The proof of Theorem 6.2 is in Appendix C.
-In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures.
-Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix H.
+This theoretical result provides a precise “invariance lever”: it allows us to arbitrarily inflate parameter norms while leaving the learned function $f(\theta)$—and thus the test error—mathematically identical at every step.
+This mechanism serves as a crucial control for our audit: if a generalization measure is sensitive to parameter magnitude (like path norm or standard PAC-Bayes), it will explode under this transformation, even though generalization performance remains identical.
+The formal definitions, the statement of Theorem C.2, and its proof are provided in Appendix C. Empirical results utilizing this lever are detailed in Appendix H.
 8 Quantitative fragility: a scale‑free instability score
 The heuristic stress tests in Sections 35 show that several popular post‑mortem measures are fragile.
-Here we provide a complementary quantitative lens: scale‑free statistics that capture how much a measure fluctuates among runs with (nearly) the same test error.
-These quantitative metrics of fragility can facilitate larger-scale comparisons between measures.
-Setup.
-Fix a dataset/architecture group $g$ with trained runs $\mathcal{S}_{g}$.
-For run $r\in\mathcal{S}_{g}$, let $\widehat{\varepsilon}_{r}\in[0,1]$ denote test error and $C_{r}\in(0,\infty)$ the value of a post‑mortem measure $C$ (e.g., a norm proxy).
-For an absolute error tolerance $\delta>0$, define the close‑error set
-$\Pi_{\delta}(\mathcal{S}_{g})\;=\;\bigl\{(r,s)\in\mathcal{S}_{g}^{2}:\ r<s,\ \lvert\widehat{\varepsilon}_{r}-\widehat{\varepsilon}_{s}\rvert\leq\delta\bigr\}.$ (12)
-Definition 7.1 (Conditional Measure Spread (CMS)).
-The CMS of $C$ in group $g$ at tolerance $\delta$ is
-$\mathrm{CMS}_{\delta}(C;g)\;=\;\operatorname{median}_{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g})}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (13)
-Interpretation.
-$e^{\mathrm{CMS}_{\delta}}$ is the typical multiplicative factor by which $C$ varies across runs that are indistinguishable in test error up to $\delta$ (e.g., $\mathrm{CMS}_{\delta}\!\approx\!\ln 2$ means a factor‑two swing).
-Randomness adjustment.
-To separate hyperparameter induced jumpiness from random‑seed jitter, let $H_{r}$ collect the training hyperparameters for run $r$ (e.g., optimizer, learning rate, schedule flags, early‑stopping rule, training‑set size), and let $\textsf{seed}_{r}$ be its random seed.
-Split close‑error pairs into
-$\Pi_{\delta}^{\text{seed}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}=H_{s},\ \textsf{seed}_{r}\neq\textsf{seed}_{s}\bigr\},$ (14)
-$\Pi_{\delta}^{\text{inter}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}\neq H_{s}\bigr\}.$ (15)
-Define the corresponding spreads
-$\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{seed}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$, (16)
-$\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{inter}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (17)
-Definition 7.2 (eCMS (Excess CMS)).
-The seed‑adjusted excess at matched error is
-$\mathrm{eCMS}_{\delta}(C;g)\;=\;\Bigl[\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)-\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)\Bigr]_{+}.$ (18)
-Aggregation across groups.
-Let $\mathcal{G}$ be the set of dataset/architecture groups. We summarize by medians across groups (and report coverage):
-$\mathrm{CMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{CMS}_{\delta}(C;g)\bigr\},$ (19)
-$\mathrm{eCMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{eCMS}_{\delta}(C;g)\bigr\}.$ (20)
-This “median‑of‑medians” is robust to heterogeneous group sizes and outliers.
-How to read the scores.
-  • CMS. Values near 0 indicate a stable measure among equal‑error runs; larger values mean greater local instability. (For reference, $\ln 2\approx 0.69$, $\ln 3\approx 1.10$ in the same log units.)
-  • eCMS. Values near 0 indicate that most equal‑error variability is seed‑level; large eCMS points to hyperparameter‑sensitive jumpiness beyond seeds.
-  • Effect of $\delta$. Decreasing $\delta$ tightens the notion of “equal‑error.” Scores that remain large as $\delta\!\downarrow\!0$ reflect genuinely volatile measures.
-Properties and protocol.
-Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$).
-Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (12)–(15) and the fraction of groups covered.
-Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
-Full experimental details, formal algorithms and tabulated results can be found in Appendix E.
-8 Conclusion and discussion
-We take a pragmatic view of generalization bounds: in modern deep learning they function best as measures to be judged by how they behave across data and training regimes rather than by worst‑case tightness.
-Read this way, fragility becomes a key property to audit.
-With the task fixed and accuracy essentially steady, small and reasonable pipeline tweaks—changing the learning rate by a factor of ten, swapping Adam for momentum SGD, toggling early stopping—can bend surrogate curves, invert trends, or inflate their scale.
-When the task complexity changes, via label corruption or by moving from MNIST to FashionMNIST to CIFAR‑10, several surrogates fail to preserve the expected ordering or slope.
-Our scale‑invariance construction makes this failure mode explicit: one can inflate parameter magnitudes without changing the predictor, and magnitude‑sensitive post‑mortems then balloon while test error does not.
-This gap underscores that properties of a single parameter vector need not reflect properties of the learned function.
-The comparison between post‑mortem diagnostics and the marginal‑likelihood PAC–Bayes route in function space sharpens the lesson.
-This predictor can capture key trends with data complexity, including the slopes of learning curves.
-Given that the interaction between data and architecture is one of the key components of DNN performance that needs explaining, the success of this bound, and its simple interpretation in terms of inductive bias at initialisation, provides a concrete foil against which parameter‑space surrogates can be refined.
-Looking forward, several concrete directions could strengthen this position:
-  • Design reparameterization‑invariant, data‑aware post‑training diagnostics that borrow the Occam flavour of marginal likelihood while remaining practical at scale;
-  • build a public fragility benchmark and automated audit harness covering optimizer/schedule swaps, scale‑symmetry probes, and data‑complexity sweeps.
-  • Develop approximate function‑space surrogates (finite‑width corrections, amortized evidence estimators, ensembles as posteriors over functions) to make invariance‑friendly predictors usable in routine training.
-  • Tie diagnostics to training dynamics by discouraging scale‑only drift after interpolation or by incorporating function‑space objectives during training.
-  • Extend evaluation beyond vision classification to sequence models, generative modelling, reinforcement learning, and explicit distribution shift, testing whether the same invariance principles hold.
+To facilitate large-scale comparisons, we developed a complementary quantitative framework that assigns a specific instability score to a measure.
+We introduce two statistics: Conditional Measure Spread (CMS), which captures the fluctuation of a measure across runs with nearly identical test error, and Excess CMS (eCMS), which isolates the component of that fluctuation driven specifically by hyperparameter changes (net of random seed variance).
+Summary of results.
+As detailed in Appendix E, we computed these scores for a wide range of measures.
+We found that parameter-count proxies are stable (eCMS $\approx 0$), while path norms, spectral distances, and standard PAC-Bayes surrogates exhibit high instability (large eCMS), quantitatively confirming the qualitative fragility observed in our stress tests.
+See Appendix E for the formal definitions (Definitions E.1 and E.2) and the full tabulated results.
 9 Alternative views
 Reasonable researchers can disagree with our emphasis on parameter-space fragility. We highlight several viable counterpositions and address them in turn:
-  • Optimizer sensitivity is signal, not noise. Because implicit bias drives which predictor is learned, a useful measure should move with the training pipeline, not be invariant to it.
-  • Post‑mortem measures can be stabilized.
-With the right reparameterization‑aware definitions (e.g., normalized sharpness, margin distributions, distance from initialization), much of the apparent fragility disappears.
-  • Function‑space GP baselines are mis‑specified for modern, finite‑width practice.
-GP priors, absence of data augmentation in the prior, and finite‑width effects limit how decisively marginal‑likelihood predictors should be used as a reference.
 Optimizer sensitivity as a feature.
 It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalization measures.
-This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019).
+This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (soudry2018the; smith2017donb; simsekli2019tail).
 From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
@@
 A second objection holds that many post‑training measures already address scale and parameterization issues.
-Normalized or reparameterization‑aware sharpness correlates more reliably with generalization than raw curvature (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022); modeling the distribution of margins rather than the minimum improves predictiveness across runs (Jiang et al., 2018); reference‑aware surrogates such as distance from initialization or movement from pretraining reduce width and scale artifacts and come with supporting analyses (Li et al., 2018; Neyshabur et al., 2017a; Zhou et al., 2021).
+Normalized or reparameterization‑aware sharpness correlates more reliably with generalization than raw curvature (dinh2017sharpb; tsuzuku2020normalized; kim2022scale; jang2022reparametrization); modeling the distribution of margins rather than the minimum improves predictiveness across runs (jiang2018predicting); reference‑aware surrogates such as distance from initialization or movement from pretraining reduce width and scale artifacts and come with supporting analyses (li2018on; neyshabur2017exploring; zhou2021understanding).
 On this view, fragility largely reflects naive implementations, not intrinsic flaws.
@@
 Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination.
-While GPs are known to closely match DNN performance on the datasets we test in this paper (Lee et al., 2020b), it is hard to see how one could untangle the effects of hyperparameters or optimiser choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address.
+While GPs are known to closely match DNN performance on the datasets we test in this paper (lee2019wide), it is hard to see how one could untangle the effects of hyperparameters or optimizer choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address.
 More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified.
 In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys.
-Therefore, the GP route’s stability could stem from omitting factors that matter in practice (Lee et al., 2020b; Valle-Pérez & Louis, 2020).
+Therefore, the GP route’s stability could stem from omitting factors that matter in practice (lee2019wide; valleprez2020generalization).
 Response.
 We agree that one of the most important questions that needs explaining in DNNs is how, when, and why finite‑width networks outperform their infinite-width GP limits, in particular for settings where feature-learning is important (see e.g.
-(Goring et al., 2025) and references therein).
+(goring2025feature) and references therein).
 We use the GP based marginal‑likelihood bound as a calibration tool, not an oracle: it naturally encodes desirable invariances (insensitivity to parameter scale and optimizer path once the dataset is fixed) and for the cases we study, consistently tracks data difficulty and learning‑curve scaling across families.
 Those properties make it a valuable foil for stress‑testing post‑mortem bounds.
-Where the GP prior is clearly misspecified (e.g., heavy augmentation or domains where feature-leaning is key), our recommendation is empirical: rerun the same fragility audit.
+Where the GP prior is clearly misspecified (e.g., heavy augmentation or domains where feature-learning is key), our recommendation is empirical: rerun the same fragility audit.
 If the GP predictor ceases to track error while a reparameterization‑invariant post‑mortem does, that is evidence for the importance of a post‑mortem analysis in that regime.
-Synthesis.
-These counterarguments motivate a middle path.
-Rather than elevating any single diagnostic, we advocate (i) reporting multiple invariance‑aware post‑mortem measures, (ii) auditing them with our stressors (learning‑curve shape, temporal drift after interpolation, and data‑complexity response), and (iii) anchoring interpretation with at least one function–space reference when feasible.
-This recognizes optimizer‑dependent mechanisms while discouraging conclusions drawn from features (e.g., raw scale) that can be changed without affecting predictions.
-Impact Statement
-This is a position paper about evaluation practice.
-The main impact is methodological: a simple audit protocol and reporting guidelines that encourage robustness checks for generalization measures.
-Potential positive impacts include more reliable model selection and clearer communication of limitations; potential negative impacts are minimal, as we do not propose deployable systems or datasets.
+10 Call to action: a better standard for evaluation
+Our position is not merely that current measures are fragile, but that the community must adopt new standards to diagnose and mitigate this fragility.
+We propose three concrete actions for authors, reviewers, and benchmark designers.
+1.
+Mandate the Fragility Audit.
+Just as error bars are expected for experimental results, a “stability check” should accompany any proposed generalization measure.
+Authors introducing a new bound or proxy should report its behavior under the Minimum Fragility Stress-Test:
+  • The Hyperparameter Sweep: Plot the measure against learning rate and optimizer changes.
+If the measure shifts by orders of magnitude while test error remains constant (as seen in Fig.
+1), this sensitivity must be reported.
+  • The Data-Complexity Check: Demonstrate that the measure correctly ranks datasets by difficulty (e.g., MNIST vs. CIFAR-10) or tracks label noise (Fig. 3).
+  • The Temporal Check: Report the measure’s post-interpolation slope. A valid generalization measure should not drift indefinitely after the training error has hit zero.
+2.
+Prioritize Invariance in Design.
+The community should pivot from deriving tighter worst-case bounds to designing invariant measures.
+We recommend developing post-mortem diagnostics that are explicitly invariant to layerwise reparameterizations and scale transformations (such as the Exp++ mechanism in §7).
+Measures based on raw parameter magnitudes should be avoided in favor of relative or normalized metrics, unless scale dependence is explicitly justified.
+3.
+Develop Function-Space Surrogates.
+Our results with the ML-PACBayes bound suggest that function-space quantities naturally avoid parameter-space fragility.
+A major open challenge is to develop computable approximations of function-space complexity—such as approximate marginal likelihood or function-space entropy—that scale to modern deep networks.
+We call for a research program dedicated to estimating these quantities efficiently, potentially bridging the gap between the robustness of GP limits and the feature-learning capabilities of finite-width networks.
+Conclusion.
+In summary, we have argued that generalization measures in deep learning must be judged not only by their tightness but also by their robustness.
+Through systematic auditing, we demonstrated that many popular post-mortem bounds—including path norms, spectral measures, and standard PAC-Bayes surrogates—are fragile: they can be made to vary wildly, trend in reverse, or collapse entirely under benign training changes that leave the underlying network’s performance effectively unchanged.
+More subtle, but important, fragilities include failing to capture basic generalisation trends with data size or data complexity.
+This fragility is likely more pronounced for measures that probe parameter space rather than function space.
+By contrast, a simple function-based marginal-likelihood PAC-Bayes bound, which captures the inductive bias of the architecture, tracks data complexity and learning-curve scaling, though it cannot distinguish optimizer effects.
+Its performance provides a natural benchmark for future, more sophisticated measures that also capture post-training effects.
+We contend that the current practice of evaluating post-training diagnostics—often limited to mild variations in training pipelines—is insufficient to expose true fragility.
+By adopting the more rigorous audits and invariance principles proposed in this work, the field can move toward measures that genuinely reflect and, more importantly, help us understand the mechanisms of generalization.
 Appendix A Training-hyperparameter fragility for all measures
-In this section we give concise definitions of the measures we study (aligning notation with prior large-scale evaluations) and provide additional experimental evidence beyond the main text (Dziugaite et al., 2020b; Jiang et al., 2019b).
+In this section we give concise definitions of the measures we study (aligning notation with prior large-scale evaluations) and provide additional experimental evidence beyond the main text (dziugaite2020in; jiang2019fantastic).
 We then present figure-backed comparisons where small learning-rate or optimizer tweaks trigger qualitatively different curve shapes for the surrogate—plateaus, rebounds, late spikes, and crossings—while the accompanying accuracy curves remain comparatively calm.
 A.1 Measures considered
-We adopt the categories and normalizations used by prior experimental studies of generalization measures (Jiang et al., 2019b; Dziugaite et al., 2020b); for exact constants and implementation choices, see App.
-C.6 of Dziugaite et al.
-(2020b).
+We adopt the categories and normalizations used by prior experimental studies of generalization measures (jiang2019fantastic; dziugaite2020in); for exact constants and implementation choices, see App.
+C.6 of dziugaite2020in.
 Let a $d$-layer network have layer weights $\{W_{i}\}_{i=1}^{d}$ with initialization $\{W_{i}^{0}\}$; write $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ for Frobenius and spectral norms; let $n$ denote training-set size and $\gamma$ a robust (e.g., 10th-percentile) training margin.
@@
 We follow the combined ratio used in prior large‑scale studies (App.
-C.6 of Dziugaite et al., 2020a), denoted FRO_OVER_SPEC, which normalizes Frobenius quantities by spectral ones to reduce raw scale effects.
+C.6 of dziugaite2020in), denoted FRO_OVER_SPEC, which normalizes Frobenius quantities by spectral ones to reduce raw scale effects.
 We report it alongside its constituents in our plots.
@@
 Across all families above, the figures make a consistent point: changing only the learning rate or swapping the optimizer can reshape the surrogate’s learning curve—introducing plateaus, rebounds, late spikes, or crossings—without a commensurate shift in test accuracy.
-This shape-level fragility extends far beyond Path norms and cautions against reading causal explanations of generalization from any single surrogate without dedicated stress tests (Dziugaite et al., 2020b; Jiang et al., 2019b).
+This shape-level fragility extends far beyond Path norms and cautions against reading causal explanations of generalization from any single surrogate without dedicated stress tests (dziugaite2020in; jiang2019fantastic).
 Appendix B Additional temporal behavior results: optimizer sensitivity across measure families
@@
 For temporal behavior of path norms, see Section 4, where those panels are discussed in detail.
-Appendix C Proof for theorem 6.2
-In this section we prove Theorem 6.2. We first reintroduce the scale invariance lemma from Arora et al. (2018b); Li & Arora (2019) which is the key source of intuition about scale-invariant networks.
+Appendix C Theory of scale-invariance: definitions and proofs
+In this section, we provide the formal definition of scale-invariant networks and the theorem regarding the equivalence of schedules (the ”Exp++” factor) discussed in Section 7.
+We then provide the full proof.
+C.1 Definitions and Theorem Statement
+Definition C.1 (Scale invariant neural networks).
+Consider a parameterized neural network $f(\boldsymbol{\theta})$. We say $f$ is scale invariant if
+$\forall c\in\mathbb{R}^{+},f(\boldsymbol{\theta})=f(c\boldsymbol{\theta})$ (1)
+Theorem C.2 (Equivalence of schedules in scale-invariant nets).
+[…]
+$\Delta_{\lambda}$ $\coloneqq(1-\gamma)^{2}-2(1+\gamma)\lambda\eta_{0}+(\lambda\eta_{0})^{2},$ (2)
+$\Xi(\alpha)$ $\coloneqq\frac{\alpha^{2}-\alpha(1-\lambda\eta_{0}+\gamma)+\gamma}{\eta_{0}}.$ (3)
+Define the interval endpoints
+$\alpha_{L}$ $\coloneqq\frac{\gamma}{\,1-\lambda\eta_{0}+\gamma\,}$, (4)
+$\alpha_{-}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}-\sqrt{\Delta_{\lambda}}}{2},$ (5)
+$\alpha_{+}$ $\coloneqq\frac{1+\gamma-\lambda\eta_{0}+\sqrt{\Delta_{\lambda}}}{2},$ (6)
+and set
+$\mathcal{I}\coloneqq(\alpha_{L},\ \alpha_{-}]\ \cup\ [\alpha_{+},\,1).$ (7)
+Consider the two updates (with $\theta_{-1}=\theta_{0}$, $\tilde{\theta}_{0}=\theta_{0}$, $\tilde{\theta}_{-1}=\alpha\,\theta_{-1}$):
+$\frac{\theta_{t}-\theta_{t-1}}{\eta_{0}}$ $=\frac{\gamma(\theta_{t-1}-\theta_{t-2})}{\eta_{0}}-\nabla_{\theta}\!\Big(L(\theta_{t-1})+\tfrac{\lambda}{2}\|\theta_{t-1}\|_{2}^{2}\Big),$ (A)
+$\frac{\tilde{\theta}_{t}-\tilde{\theta}_{t-1}}{\tilde{\eta}_{t}}$ $=\frac{\gamma(\tilde{\theta}_{t-1}-\tilde{\theta}_{t-2})}{\tilde{\eta}_{t-1}}-\nabla_{\theta}\!\Big(L(\tilde{\theta}_{t-1})+\tfrac{\tilde{\lambda}_{t}}{2}\|\tilde{\theta}_{t-1}\|_{2}^{2}\Big).$ (B)
+If $\alpha\in\mathcal{I}$, then the schedules
+$\tilde{\eta}_{t}$ $=\eta_{0}\,\alpha^{-2t-1},$ (8)
+$\tilde{\lambda}_{t}$ $=\Xi(\alpha)\,\alpha^{2t-1}+\frac{\gamma(1-\alpha)}{\eta_{0}\,\alpha}\,\mathbbm{1}\{t=0\}$ (9)
+ensure $\tilde{\theta}_{t}=\alpha^{-t}\theta_{t}$ for all $t\geq 0$; hence (A) and (B) generate identical functions.
+Remark C.3 (Parameter range and nonnegativity).
+Assume
+$\lambda\eta_{0}$ $\leq(1-\sqrt{\gamma})^{2},$ (10)
+$\alpha_{L}$ $<\alpha_{-}\quad\text{(equivalently, $\ \Delta_{\lambda}\geq 0$)}.$ (11)
+Then the interval $\mathcal{I}$ above is nonempty and the square‑root term is real; both conditions hold for common hyperparameters in practice.
+C.2 Proof of Theorem C.2
+We first reintroduce the scale invariance lemma from arora2018theoretical; li2019an which is the key source of intuition about scale-invariant networks.
 Lemma C.4 (Scale-invariant networks).
@@
 $\boldsymbol{\theta}$ and set $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$, we immediately arrive at the result.
-We give the technical definition of the commonly used training algorithm SGD with momentum and weight decay (WD) (with respect to the L2 norm of the parameters) using a convenient form given in Li & Arora (2019), which is equivalent to the default implementation in Pytorch (Paszke et al., 2017).
+We give the technical definition of the commonly used training algorithm SGD with momentum and weight decay (WD) (with respect to the L2 norm of the parameters) using a convenient form given in li2019an, which is equivalent to the default implementation in Pytorch (paszke2017automatic).
 Definition C.5 (SGD with momentum and WD).
@@
 $=$ $\text{GD}_{t}^{\rho}\left(\boldsymbol{\theta},\eta,\boldsymbol{\theta}^{\prime},\eta\right)$
-The range of $\alpha$ can be easily shown by combining the following two constraints and assuming remark 6.3 is true:
+The range of $\alpha$ can be easily shown by combining the following two constraints and assuming remark C.3 is true:
   • $\alpha\in\left(0,1\right];$
   • $\frac{(\rho+\gamma)}{\alpha}-\frac{\gamma}{\alpha^{2}}\in\left(0,1\right]$
-Now we are ready to prove theorem 6.2.
-Proof of Theorem 6.2.
+Now we are ready to prove theorem C.2.
+Proof of Theorem C.2.
 From the assumption we have the following equivalence between the boundary conditions of the two series of states:
@@
 All panels use 10,000 training samples.
-Appendix E Quantitative fragility: detailed results
-This appendix extends Section 7.
+Appendix E Quantitative fragility: detailed definitions and results
+This appendix provides the formal definitions, algorithmic details, and full results for the quantitative analysis summarized in Section 8.
+E.1 Formal Definitions
+To quantify fragility, we define metrics that capture how much a generalization measure fluctuates among runs with nearly identical test error.
+Setup.
+Fix a dataset/architecture group $g$ with trained runs $\mathcal{S}_{g}$.
+For run $r\in\mathcal{S}_{g}$, let $\widehat{\varepsilon}_{r}\in[0,1]$ denote test error and $C_{r}\in(0,\infty)$ the value of a post‑mortem measure $C$ (e.g., a norm proxy).
+For an absolute error tolerance $\delta>0$, define the close‑error set
+$\Pi_{\delta}(\mathcal{S}_{g})\;=\;\bigl\{(r,s)\in\mathcal{S}_{g}^{2}:\ r<s,\ \lvert\widehat{\varepsilon}_{r}-\widehat{\varepsilon}_{s}\rvert\leq\delta\bigr\}.$ (15)
+Definition E.1 (Conditional Measure Spread (CMS)).
+The CMS of $C$ in group $g$ at tolerance $\delta$ is
+$\mathrm{CMS}_{\delta}(C;g)\;=\;\operatorname{median}_{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g})}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (16)
+Interpretation.
+$e^{\mathrm{CMS}_{\delta}}$ is the typical multiplicative factor by which $C$ varies across runs that are indistinguishable in test error up to $\delta$ (e.g., $\mathrm{CMS}_{\delta}\!\approx\!\ln 2$ means a factor‑two swing).
+Randomness adjustment.
+To separate hyperparameter induced jumpiness from random‑seed jitter, let $H_{r}$ collect the training hyperparameters for run $r$ (e.g., optimizer, learning rate, schedule flags, early‑stopping rule, training‑set size), and let $\textsf{seed}_{r}$ be its random seed.
+Split close‑error pairs into
+$\Pi_{\delta}^{\text{seed}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}=H_{s},\ \textsf{seed}_{r}\neq\textsf{seed}_{s}\bigr\},$ (17)
+$\Pi_{\delta}^{\text{inter}}(g)$ $=\bigl\{(r,s)\in\Pi_{\delta}(\mathcal{S}_{g}):\ H_{r}\neq H_{s}\bigr\}.$ (18)
+Define the corresponding spreads
+$\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{seed}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$, (19)
+$\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)$ $=\operatorname{median}_{(r,s)\in\Pi_{\delta}^{\text{inter}}(g)}\bigl\lvert\log C_{r}-\log C_{s}\bigr\rvert$. (20)
+Definition E.2 (eCMS (Excess CMS)).
+The seed‑adjusted excess at matched error is
+$\mathrm{eCMS}_{\delta}(C;g)\;=\;\Bigl[\mathrm{CMS}^{\text{inter}}_{\delta}(C;g)-\mathrm{CMS}^{\text{seed}}_{\delta}(C;g)\Bigr]_{+}.$ (21)
+Aggregation across groups.
+Let $\mathcal{G}$ be the set of dataset/architecture groups. We summarize by medians across groups (and report coverage):
+$\mathrm{CMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{CMS}_{\delta}(C;g)\bigr\},$ (22)
+$\mathrm{eCMS}_{\delta}^{\mathrm{med}}(C)$ $=\operatorname{median}_{g\in\mathcal{G}}\bigl\{\mathrm{eCMS}_{\delta}(C;g)\bigr\}.$ (23)
+This “median‑of‑medians” is robust to heterogeneous group sizes and outliers.
+How to read the scores.
+  • CMS. Values near 0 indicate a stable measure among equal‑error runs; larger values mean greater local instability. (For reference, $\ln 2\approx 0.69$, $\ln 3\approx 1.10$ in the same log units.)
+  • eCMS. Values near 0 indicate that most equal‑error variability is seed‑level; large eCMS points to hyperparameter‑sensitive jumpiness beyond seeds.
+  • Effect of $\delta$. Decreasing $\delta$ tightens the notion of “equal‑error.” Scores that remain large as $\delta\!\downarrow\!0$ reflect genuinely volatile measures.
+Properties and protocol.
+Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$).
+Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (15)–(18) and the fraction of groups covered.
+Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
+E.2 Experimental Details
 We first summarize the experimental setup, and then give concise, implementable procedures for our two instability scores: CMS and its seed‑adjusted variant eCMS.
 We close with a brief reading guide that highlights what can be learned from the summary tables at Table 1 ($\delta=0.01$), Table 2 ($\delta=0.02$), and Table 3 ($\delta=0.05$).
-For full definitions of the measures listed in those tables, see (Dziugaite et al., 2020b).
+For full definitions of the measures listed in those tables, see (dziugaite2020in).
 Experimental setup.
-We study two datasets—FashionMNIST and CIFAR‑10—across three architectures: Network‑in‑Network (NiN) (Lin et al., 2013), ResNet‑50 (RN50), and DenseNet‑121 (DN121).
+We study two datasets—FashionMNIST and CIFAR‑10—across three architectures: Network‑in‑Network (NiN) (lin2013network), ResNet‑50 (RN50), and DenseNet‑121 (DN121).
 The training pipeline varies along three axes: (i) learning rate $\texttt{lr}\in\{0.001,0.0032,0.0063,0.01,0.0158,0.05,0.1\}$; (ii) optimizer $\in\{\textsc{Adam},\textsc{SGD}+\textsc{momentum}\}$; and (iii) stopping criterion: either the first epoch reaching $100\%$ training accuracy, or the first epoch whose training cross‑entropy drops below $0.01$.
@@
 Appendix F Post-mortem vs. ML-PACBayes
-We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of Valle-Pérez & Louis (2020).
+We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of valleprez2020generalization.
 The bound controls the test error of a hypothesis sampled from a Bayesian posterior over functions (not parameters); the key capacity term is the marginal likelihood (Bayesian evidence) of the data under the architecture’s Gaussian‑process (GP) limit.
-All figure numbers below refer to Valle-Pérez & Louis (2020).
+All figure numbers below refer to valleprez2020generalization.
 Definition F.1 (Marginal-likelihood PAC–Bayes bound).
@@
   • Data complexity.
-The marginal‑likelihood bound tracks dataset difficulty: it increases with label corruption and preserves the canonical ordering MNIST << Fashion‑MNIST << CIFAR‑10; see Figure 1 in (Valle-Pérez & Louis, 2020).
+The marginal‑likelihood bound tracks dataset difficulty: it increases with label corruption and preserves the canonical ordering MNIST << Fashion‑MNIST << CIFAR‑10; see Figure 1 in (valleprez2020generalization).
 In our audits, several post‑mortem measures—including PACBAYES_orig—either flattened under corruption or were not able to catch the ordering, even when the test error did.
   • Learning‑curve scaling.
-The marginal‑likelihood bound tracks learning‑curve scaling: it mirrors the empirical power‑law in $n$ and clusters exponents primarily by dataset; see Figures 2–5 in (Valle-Pérez & Louis, 2020).
+The marginal‑likelihood bound tracks learning‑curve scaling: it mirrors the empirical power‑law in $n$ and clusters exponents primarily by dataset; see Figures 2–5 in (valleprez2020generalization).
 By contrast, some post‑mortem measures barely moved with $n$; others bent the wrong way or mixed optimizer effects with data effects (cf.
-(Nagarajan & Kolter, 2019)).
+(nagarajan2019uniform)).
   • Temporal behavior (overtraining invariance).
@@
 In our temporal‑behavior experiments, several norm‑based post‑mortem measures (e.g., $\ell_{2}$/spectral/path norms) kept growing during overtraining while the generalization error stayed roughly unchanged; the GP bound remained stable.
-(This echoes broader observations that longer training can leave test error flat or improved (Hoffer et al., 2017), and that norms can diverge under separable losses without hurting classification error (Soudry et al., 2018).)
+(This echoes broader observations that longer training can leave test error flat or improved (hoffer2017train), and that norms can diverge under separable losses without hurting classification error (soudry2018the).)
   • Training‑pipeline changes.
-Many post‑mortem measures swing with optimizer, batch size, explicit/implicit regularization, and early stopping—sometimes tracking curvature artifacts rather than out‑of‑sample error (e.g., Keskar et al., 2017; Hochreiter & Schmidhuber, 1997; Jiang et al., 2019b; Dziugaite et al., 2020b).
+Many post‑mortem measures swing with optimizer, batch size, explicit/implicit regularization, and early stopping—sometimes tracking curvature artifacts rather than out‑of‑sample error (e.g., keskar2016on; hochreiter1997flatb; jiang2019fantastic; dziugaite2020in).
 The GP marginal‑likelihood bound is, by construction, insensitive to these knobs: it depends on the architecture’s function prior and the data, not on the path taken through parameter space.
@@
 First, the bound evaluates an Occam factor in function space: architectures that place high prior mass on data‑compatible functions earn better evidence, naturally capturing dataset ordering and learning‑curve slopes.
-Second, there is a credible external yardstick: infinite‑width GP predictors (NNGP/NTK) often quantitatively predict finite‑width DNN generalization trends (Lee et al., 2020a).
+Second, there is a credible external yardstick: infinite‑width GP predictors (NNGP/NTK) often quantitatively predict finite‑width DNN generalization trends (lee2020finite).
 Together, these explain why Figures 1–5 show greater stability for the GP‑based approach than we observe for post‑hoc, parameter‑space measures.
@@
 Post‑mortem scores probe properties of a single trained parameter vector (curvature, norms, compressibility), entangling optimization details and reparameterization choices with generalization.
-That enterprise is important—and the community has argued forcefully that such post‑training diagnostics deserve attention (e.g., compression and robustness perspectives in Arora et al., 2018a, 2019; Jiang et al., 2019b; Dziugaite et al., 2020b)—but our evidence suggests their predictions are fragile under routine perturbations.
+That enterprise is important—and the community has argued forcefully that such post‑training diagnostics deserve attention (e.g., compression and robustness perspectives in arora2018strongerc; arora2019fine; jiang2019fantastic; dziugaite2020in)—but our evidence suggests their predictions are fragile under routine perturbations.
 By contrast, the GP bound targets the distribution over functions implied by architecture and data.
@@
 We instantiate the symmetry in a fully connected, scale‑invariant network (normalization after each hidden linear layer; the final linear readout is frozen), trained on MNIST with SGD + momentum.
-We sweep a single control, the Exp++ multiplicative factor $\alpha$ from Theorem 6.2, which makes the learning rate grow exponentially across steps, and repeat the sweep both without and with weight decay.
+We sweep a single control, the Exp++ multiplicative factor $\alpha$ from Theorem C.2, which makes the learning rate grow exponentially across steps, and repeat the sweep both without and with weight decay.
 For each run we log parameter‑space PAC–Bayes proxies (including magnitude‑aware variants), path‑norm proxies (with/without margin normalization), and test error.
@@
 Together, the panels illustrate a clean separation between parameter scale (which Exp++ manipulates) and predictive behavior (which stays essentially fixed).
+Appendix I Additional Compression Bounds Results
+In Section 6, we analyzed the breakdown of compression bounds for CIFAR-10 (Intrinsic Dimension 3500) and MNIST.
+Here, we provide the corresponding omitted results for CIFAR-10 with a higher intrinsic dimension ($ID=10000$).
+(a) CIFAR-10 (ID 10000) Bounds
+(b) CIFAR-10 (ID 10000) Raw KL
+Figure S20: Compression Bounds with Higher Intrinsic Dimension (CIFAR-10).
+These plots correspond to the setting where the intrinsic dimension is set to 10,000.
+Left: The bound components versus sample size.
+Right: The raw KL divergence.
+Consistent with the main text analysis, the complexity term (KL) remains flat despite the increase in sample size.